### v6

# The local converse theorem for $\widetilde{\mathrm{Sp}}_{2n}$

Abstract.
In this paper, we establish the local converse theorem and the stability of local gamma factors for $\widetilde{\mathrm{Sp}}_{2n}$ via the precise local theta correspondence between $\widetilde{\mathrm{Sp}}_{2n}$ and $\mathrm{SO}_{2n+1}$ over local fields of characteristic not equal to 2. As an application, we can prove the rigidity theorem for irreducible generic cuspidal automorphic representations of $\widetilde{\mathrm{Sp}}_{2n}$ over number fields.
Key words and phrases:
gamma factors, metaplectic groups, local converse theorem, theta correspondence
2020 Mathematics Subject Classification:
11F70, 22E50
1. Introduction
Let $F$ be a non-archimedean local field of characteristic not equal to 2 (i.e., a finite extension of $\mathbb{Q}_{p}$ for any odd prime $p$ or $\mathbb{F}_{q}((t))$ for $q=p^{k}$), and let $W$ be a $2n$-dimensional symplectic space over $F$. Define $\widetilde{\mathrm{J}_{n}}=\widetilde{\mathrm{Sp}}(W)$ as the metaplectic group of $W$, which is the double covering group of $\mathrm{J}_{n}=\mathrm{Sp}(W)$, the isometry group of $W$.
We introduce $\psi$, a non-trivial additive character of $F$. For $\lambda\in F^{\times}$, we define the character $\psi_{\lambda}$ of $F$ as $\psi_{\lambda}(x)\coloneqq\psi(\lambda x)$. Let $\mathrm{U}^{\prime}$ be the unipotent radical of a Borel subgroup of $\mathrm{Sp}(W)$, and let $\mu_{\lambda}^{\prime}$ be a certain non-trivial character of $\mathrm{U}^{\prime}(F)$ that depends on $\psi$ and $\lambda$ (see Section 2.2 for its definition). In this paper, we prove the following local converse theorem for quasi-split metaplectic groups under the following hypothesis in $\operatorname{char}(F)=p\neq 2$ case. (There is no hypothesis in $\operatorname{char}(F)=0$ case.)
Working Hypothesis.
The $\gamma$-factors for $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$ are properly defined in $\operatorname{char}(F)=p\neq 2$ cases. Furthermore, they satisfy natural properties of $\gamma$-factors. (For the precision of the natural properties, see (i)-(vii) except (vi) in Proposition 2.1 )
Theorem 1.1 (Local Converse Theorem for $\widetilde{\mathrm{Sp}}_{2n}$).
Assume the above working hypothesis on $\operatorname{char}(F)=p\neq 2$ case. Let $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ be irreducible admissible $\mu_{\lambda}^{\prime}$-generic representations of $\widetilde{\mathrm{J}_{n}}(F)$, satisfying the condition
$\gamma(s,\widetilde{\pi}\times\sigma,\psi_{\lambda})=\gamma(s,\widetilde{\pi}^{\prime}\times\sigma,\psi_{\lambda})$
holds for all irreducible supercuspidal representation $\sigma$ of $\mathrm{GL}_{i}(F)$ with $1\leq i\leq n$. Then, it follows that
$\widetilde{\pi}\simeq\widetilde{\pi}^{\prime}.$
Here, $\gamma$ denotes the local gamma factor (see Section 2.3 for a detailed explanation). The local converse theorem (LCT) has been studied in various versions concerning different groups. We provide a brief overview of some notable works related to this theorem.
The initial investigation of these types of problems can be attributed to Henniart [He], who established a (weak) version of the theorem for general linear groups. Jiang and Soudry [JS03], building upon Henniart’s work, proved a (weak) LCT for odd special orthogonal groups. They combined the global weak functorial lift from $\mathrm{SO}_{2n+1}$ to $\mathrm{GL}_{2n}$, the local descent from $\mathrm{GL}_{2n}$ to $\widetilde{\mathrm{Sp}}_{2n}$, and the Howe lifts from $\mathrm{SO}_{2n+1}$ to $\widetilde{\mathrm{Sp}}_{2n}$. Their proof involved a global-to-local argument that employed Henniart’s result. Subsequently, D. Jiang [Jng] proposed a version of the LCT for all classical groups.
Around 2017 Chai [Cha19] and Jacquet and Liu [JL18] independently refined Henniart’s result to its optimal form. Recently, this is also achieved by P. Yan, Q. Zhang [YZ23] in a different manner.
Recently, Q. Zhang [Q18, Q19] proved the supercuspidal cases of the conjecture for $\mathrm{Sp}_{2n}$ and $\mathrm{U}_{2n+1}$ using the theory of partial Bessel functions developed by Cogdell, Shahidi, and Tsai in [CST17]. In a similar vein, B. Liu and A. Hazeltine [LH22] established the conjecture for split even special orthogonal groups, while Y. Jo [Jo] refined Jiang and Soudry’s [JS03] result and extended Zhang’s [Q18] result to the generic cases. Notably, Jo’s work addressed local fields of positive characteristics as well as characteristic zero, in contrast to other LCT results that only dealt with characteristic zero cases. These proofs were purely local and did not invoke the weak Langlands functorial lifts to general linear groups, as in the recent results of [Cha19] and [JL18]. For other classical groups, Morimoto [M] demonstrated the theorem for even unitary groups by pulling back the recent results of [Cha19] and [JL18] using global (and hence) local descent methods.
With regards to the LCT for metaplectic groups, it appears plausible to apply the local descent methods by pulling back the LCT for general linear groups ([Cha19] and [JL18]). Alternatively, the theory of partial Bessel functions [CST17] used in [Q18, Q19, LH22] and [Jo] can be employed. Both approaches would require significant effort. In this paper, we propose a new approach to prove the LCT for representation of metaplectic groups using the Local Theta Correspondence (LTC) between $\mathrm{SO}_{2n+1}$ and $\widetilde{\mathrm{Sp}}_{2n}$. The LTC allows for the transfer of problems between different groups, similar to how Gan and Ichino used it to prove the local Gan-Gross-Prasad conjecture for skew-hermitian unitary groups ([GI2]) and the endoscopic classification of automorphic representations metaplectic groups ([GI3].) Inspired by these works, we aim to transfer the LCT for $\mathrm{SO}_{2n+1}$ to $\widetilde{\mathrm{Sp}}_{2n}$ using the LTC.
To successfully transfer the LCT, it is necessary to demonstrate that the LTC preserves certain properties and invariants of representations. Specifically, we consider genericity (Theorem 3.3) and $\gamma$-factors (Theorem 3.7). Indeed, these results can be found in [GS, Theorem 9.1] and [GS, Proposition 11.1, Corollary 11.2], respectively.
However, unfortunately, the proof of [GS, Theorem 9.1] is absent in the cited reference. To address this gap, we provide a complete proof demonstrating that the missing argument holds even in the case when $\operatorname{char}(F)=p$. Additionally, [GS, Proposition 11.1] relies on the general $\gamma$-factor theory, which is currently unavailable. Consequently, we furnish a comprehensive proof for this proposition using only the established theory on generic $\gamma$-factor (see Remark 3.8).
Theorem 3.3 and Theorem 3.7 form the main part of the paper and are discussed in Section 3, along with Section 4, which computes the twisted Jacquet module of the Weil representation—a crucial step in proving the preservation of genericity.
According to our knowledge, this paper represents the first use of the LTC to prove the LCT within the context of classical and metaplectic groups. This approach demonstrates the use of the LTC in transferring the LCT between different groups. Indeed, a similar method can be employed to extend the recent result on the LCT for split $\mathrm{SO}_{2n}$ to quasi-split $\mathrm{SO}_{2n}$. However, due to concerns about complicating the notation and subtle differences in the proof, we chose to address this extension separately in [HKK23].
It’s worth noting that in the case where the local field $F$ has characteristic 0, the main theorem (Theorem 1.1) can be readily derived by using Gan-Ichino and Arthur’s results concerning the local Langlands correspondence for $\widetilde{\mathrm{Sp}}_{2n}$ ([GI3], [Ar13]), in conjunction with the LCT for $\mathrm{GL}_{2n}$ and the uniqueness of generic members in $L$-packets ([At]). However, we stress that our approach is entirely independent of Arthur’s results. We believe that a proof not relying on Arthur’s results holds its intrinsic value.
In addition to the main theorem (Theorem 1.1), the paper also proves the following stability of local gamma factors for $\widetilde{\mathrm{Sp}}_{2n}$ as a byproduct of the proof.
Theorem 1.2 (Stability of local $\gamma$-factors for $\widetilde{\mathrm{Sp}}_{2n}$).
Let $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ be irreducible admissible $\mu_{\lambda}^{\prime}$-generic representations of $\widetilde{\mathrm{Sp}}_{2n}(F)$. Then there exists $l=l(\widetilde{\pi},\widetilde{\pi}^{\prime})$ such that for any quasi-character $\chi$ of $F^{\times}$, with conductor $\text{cond}(\chi)>l$, we have
$\gamma(s,\widetilde{\pi}\times\chi,\psi_{\lambda})=\gamma(s,\widetilde{\pi}^{\prime}\times\chi,\psi_{\lambda}).$
When the residual characteristic of the $p$-adic field $F$ is odd, Zhang ([Q17]) directly proved this result under the condition that $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ share the same central characters. Our result covers cases where the residual characteristic of $F$ is 2 without requiring the assumption of identical central characters. Alternatively, we can remove the assumption on the same central characters from the local functional equation of doubling zeta integrals of metaplectic groups (see [GS, Section 4.3]).
The paper also presents an application of the LCT and the stability of local gamma factors for $\widetilde{\mathrm{Sp}}_{2n}$ by proving the rigidity theorem for $\widetilde{\mathrm{Sp}}_{2n}$ over a global number field.
Theorem 1.3 (Rigidity theorem for $\widetilde{\mathrm{Sp}}_{2n}$).
Let $K$ be a global field and $\mathbb{A}$ its adele ring. Let $\mathfrak{U}^{\prime}$ be a non-trivial generic character of $\mathrm{U}^{\prime}(F)\backslash\mathrm{U}^{\prime}(\mathbb{A})$ and $\widetilde{\pi}=\otimes_{v}\widetilde{\pi_{v}}$ and $\widetilde{\pi}^{\prime}=\otimes_{v}\widetilde{\pi_{v}}^{\prime}$ be irreducible cuspidal $\mathfrak{U}^{\prime}$-generic automorphic representations of $\widetilde{\mathrm{Sp}}_{2n}(\mathbb{A})$. If $\widetilde{\pi_{v}}\simeq\widetilde{\pi_{v}}^{\prime}$ for almost all $v$, then $\widetilde{\pi}\simeq\widetilde{\pi}^{\prime}.$
The rigidity theorem states that if two irreducible cuspidal automorphic representations $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ of $\widetilde{\mathrm{Sp}}_{2n}(\mathbb{A})$ are equivalent at almost all places, then $\widetilde{\pi}$ is equivalent to $\widetilde{\pi}^{\prime}$.
The paper is organized as follows:
In Section 2, we provide the necessary background and set up for the precise formulation of the theorems. We also reduce Theorem 1.1 to its weak version, which is stated as Theorem 2.5. In Section 3, we explain the LTC and review some of its known properties. We then proceed to prove that the LTC preserves genericity and gamma factors between $\widetilde{\mathrm{Sp}}_{2n}$ and $\mathrm{SO}_{2n+1}$. This is a key component of the results presented in the paper. In Section 4, we compute the twisted Jacquet module of the Weil representation, which plays a crucial role in proving the preservation of genericity in Section 3. In Section 5, we provide a proof of Theorem 1.1 and Theorem 1.2. In Section 6, we prove the rigidity theorem for $\widetilde{\mathrm{Sp}}_{2n}$ as an application of the theorems established earlier.
1.1. Notations
  • $F$ : a non-archimedean local field of characteristic different from 2
  • $\operatorname{char}(F)$ : the characteristic of $F$
  • $q$ : the order of the residue field of $F$ such that $q=p^{e}$ for some odd prime $p$
  • $\psi$ : a non-trivial additive character of $F$
  • $\varpi$ : a uniformizer of $F$
  • $|\cdot|_{F}$ : a normalized absolute value in $F$ such that $|\varpi|_{F}=q^{-1}$
  • $\operatorname{Irr}(G)$ : the set of isomorphism classes of irreducible representations of $G$, where $G:=\mathrm{G}(F)$ is a $F$-points of a reductive group $\mathrm{G}$ defined over $F$.
  • $\operatorname{Irr}_{\operatorname{temp}}(G)$ : the set of isomorphism classes of irreducible tempered representations of $G$
  • $\operatorname{Irr}_{\operatorname{supc}}(G)$ : the set of isomorphism classes of irreducible supercuspidal representations of $G$
  • $\operatorname{Irr}_{\operatorname{unit}}(G)$ : the set of isomorphism classes of irreducible unitary representations of $G$
  • $\pi^{\vee}$ : the contragredient representation of $\pi\in\operatorname{Irr}(G)$
  • $V_{m}$ : a $12m+$ 1-dimensional orthogonal space, i.e., a $12m+$ 1-dimensional vector space over $F$ equipped with a non-degenerate symmetric bilinear form
  • $\mathrm{H}_{m}:=\mathrm{SO}(V_{m})$ : the special orthogonal group of $V_{m}$ a non-degenerate symmetric bilinear form $(\ ,\ )_{V_{m}}$
  • $W_{n}$ : a $2n$-dimensional symplectic space over $F$, i.e., a $2n$-dimensional vector space over $F$ equipped with a non-degenerate symplectic bilinear form $\langle\ ,\ \rangle_{W_{n}}$
  • $\mathrm{J}_{n}:=\mathrm{Sp}(W_{n})$ : the symplectic group of $W_{n}$
  • $\widetilde{\mathrm{J}_{n}}:=\widetilde{\mathrm{Sp}}(W_{n})$ : the metaplectic group of $W_{n}$
  • $\mathbb{I}$ : the trivial representation
  • $\operatorname{ind}_{B}^{G}$ : the unnormalized compactly supported induction for an algebraic group $G$ and its closed subgroup $B$
  • $\mathcal{S}(X)$ : the space of locally constant function of compact support on a set $X$
2. Preliminaries
In this section, we prepare the basic setup to state our main theorem. While many of the theorems and propositions we quote in this paper only address the case where $\operatorname{char}(F)=0$, most of the proofs apply equally to the positive characteristic cases. Therefore, instead of repeating the same proof in the references for the case $\operatorname{char}(F)=0$, we adopt the position that they also cover the case $\operatorname{char}(F)=p$ when it is applicable. However, when the proof of a proposition or a theorem for $\operatorname{char}(F)=0$ is not clearly stated in the references, we provide a proof that covers both cases. Furtherore, when the proofs for $\operatorname{char}(F)=0$ and $\operatorname{char}(F)=p$ are a bit different, we mention such differences in the proof.
2.1. Metaplectic and orthogonal groups
2.1.1. Symplectic groups
Let $W_{n}$ be a $2n$-dimensional vector space over the field $F$ equipped with a non-degenerate symplectic form denoted by $\langle\cdot,\cdot\rangle_{W_{n}}$. We denote the associated isometric group as $\mathrm{Sp}(W_{n})$.
We choose a specific basis $\{f_{1},\cdots,f_{n},f_{n}^{*},\cdots,f_{1}^{*}\}$ for $W_{n}$ such that the following relations hold:
$\langle f_{i},f_{j}\rangle_{W_{n}}=\langle f_{i}^{*},f_{j}^{*}\rangle_{W_{n}}=0,\quad\langle f_{i},f_{j}^{*}\rangle_{W_{n}}=\delta_{ij},$
where $\delta_{ij}$ is the Kronecker delta.
For $1\leq k\leq n$, we define
$X_{k}=\text{Span}\{f_{1},\cdots,f_{k}\}\quad\text{and}\quad X_{k}^{*}=\text{Span}\{f_{1}^{*},\cdots,f_{k}^{*}\}.$
This gives us $W_{n}=X_{k}\oplus X_{k}^{*}$.
We also define
$W_{n,k}=\text{Span}\{f_{k+1},\cdots,f_{n},f_{n}^{*},\cdots,f_{k+1}^{*}\},$
so that $W_{n}=X_{k}\oplus W_{n-k}\oplus X_{k}^{*}$.
Next, we consider the flag of isotropic subspaces
$X_{k_{1}}\subset X_{k_{1}+k_{2}}\subset\cdots\subset X_{k_{1}+\cdots+k_{r}}\subset W_{n}.$
The stabilizer of such a flag is a parabolic subgroup $\mathrm{P}^{\prime}$ of $\mathrm{Sp}(W_{n})$ whose Levi factor $\mathrm{M}^{\prime}$ is isomorphic to
$\mathrm{M}^{\prime}\simeq\mathrm{GL}_{k_{1}}\times\cdots\times\mathrm{GL}_{k_{r}}\times\mathrm{Sp}(W_{n-k_{1}-\cdots-k_{r}}),$
where each $\mathrm{GL}_{k_{i}}$ is the group of invertible linear maps on $\text{Span}\{f_{k_{i}+1},\cdots,f_{k_{i+1}}\}$.
2.1.2. Metaplectic groups
For a symplectic space $W_{n}$, metaplectic group $\widetilde{\mathrm{Sp}}(W_{n})$ is a two-fold cover of $\mathrm{Sp}(W_{n})$ defined by
$\widetilde{\mathrm{Sp}}(W_{n})=\mathrm{Sp}(W_{n})\times\{\pm{1}\}$
with group law given by
$(g_{1},\epsilon_{1})\times(g_{2},\epsilon_{2})=(g_{1}g_{2},\epsilon_{1}\epsilon_{2}\cdot c(g_{1},g_{2})),$
where $c$ is some 2-cocycle on $\mathrm{Sp}(W_{n})$ whose values lie in {±$\{\pm 1\}$. The covering group $\widetilde{\mathrm{GL}}(X_{k})$ of $\mathrm{GL}(X_{k})$ is defined by $\mathrm{GL}(X_{k})\times\{\pm 1\}$ with group law
$(g_{1},\epsilon_{1})\cdot(g_{2},\epsilon_{2})=(g_{1}g_{2},\epsilon_{1}\epsilon_{2}\cdot(\det g_{1},\det g_{2})),\quad\quad g_{1},g_{2}\in\mathrm{GL}(X_{k}),$
where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the Hilbert symbol.
For a Levi subgroup $\mathrm{M}^{\prime}\simeq\mathrm{GL}_{k_{1}}\times\cdots\times\mathrm{GL}_{k_{r}}\times\widetilde{\mathrm{Sp}}(W_{0})$ of $\mathrm{Sp}(W_{n})$, denote its double covering group $\widetilde{\mathrm{M}}^{\prime}$ by
$\widetilde{\mathrm{M}}^{\prime}\simeq\widetilde{\mathrm{GL}_{k_{1}}}\times_{\{\pm{1}\}}\cdots\times_{\{\pm{1}\}}\widetilde{\mathrm{GL}_{k_{r}}}\times_{\{\pm{1}\}}\widetilde{\mathrm{Sp}}(W_{0}).$
Any unipotent subgroup $\mathrm{N}^{\prime}$ of $\mathrm{Sp}(W_{n})$ splits uniquely in $\widetilde{\mathrm{Sp}}(W_{n})$. Therefore, we denote its splitting image in $\widetilde{\mathrm{Sp}}(W_{n})$ using the same symbol $\mathrm{N}^{\prime}$. If $\mathrm{P}^{\prime}=\mathrm{M}^{\prime}\mathrm{N}^{\prime}$ is a parabolic subgroup of $\mathrm{Sp}(W_{n})$, we call $\widetilde{\mathrm{P}}^{\prime}=\widetilde{\mathrm{M}}^{\prime}\mathrm{N}^{\prime}$ the double covering group of $\mathrm{P}^{\prime}$. Every parabolic subgroup of $\widetilde{\mathrm{Sp}}(W_{n})$ has this form.
2.1.3. Orthogonal groups
Let $V_{m}$ be a $(2m+1)$-dimensional vector space over $F$ equipped with a non-degenerate symmetric bilinear form $(\cdot,\cdot)_{V_{m}}$ of discriminant 1. There are two isomorphism classes of $V_{m}$ depending on the dimension of maximal isotropic subspaces of $V_{m}$. If it is $m$, it is denoted by $V_{m}^{+}$ and called a split quadratic space. Otherwise, the dimension of maximal isotropic subspaces of $V_{m}$ is $m-1$ and it is denoted by $V_{m}^{-}$. We call this one non-split. Define
$\epsilon(V_{m})=\begin{cases}+,\quad\text{if $V_{m}$ is split}\\ -,\quad\text{if $V_{m}$ is non-split}.\end{cases}$
Let $\mathrm{O}(V_{m})$ be the associated orthogonal group of $V_{m}$. Observe that
$\mathrm{O}(V_{m})=\mathrm{SO}(V_{m})\times\{\pm 1\},$
where $\mathrm{SO}(V_{m})$ is the special orthogonal group. Then $\mathrm{SO}(V_{m})$ is split if and only if $V_{m}$ is split. Given any irreducible admissible representation $\pi$ of $\mathrm{SO}(V_{m})$, there are exactly two extensions of $\pi$ to $\mathrm{O}(V_{m})$. It is determined by the action of $-1\in\mathrm{O}(V_{m})$; it acts on $V_{m}$ as a scalar multiple $1$ or $-1$. Therefore, the study of $\operatorname{Irr}(\mathrm{O}(V_{m}))$ is essentially that of $\operatorname{Irr}(\mathrm{SO}(V_{m}))$.
Let $\{e_{1},\cdots,e_{r}\}$ and $\{e_{1}^{*},\cdots,e_{r}^{*}\}$ be maximal isotropic subspaces of $V_{m}$ satisfying
$\langle e_{i},e_{j}\rangle_{V_{m}}=\langle e_{i}^{*},e_{j}^{*}\rangle_{V_{m}}=0,\quad\langle e_{i},e_{j}^{*}\rangle_{V_{m}}=\delta_{ij}.$
For $1\leq k\leq r$, let
$Y_{k}=\text{Span}\{e_{1},\cdots,e_{k}\}\text{ and }Y_{k}^{*}=\text{Span}\{e_{1}^{*},\cdots,e_{k}^{*}\},$
and $V_{m,k}$ be a subspace of $V_{m}$ so that $V_{m}=Y_{k}\oplus V_{m,k}\oplus Y_{k}^{*}$. If $V_{m}$ is split, let $e$ be an element in $V_{m,m}$ such that $(e,e)_{V_{m}}=1$.
Next, we consider the flag of isotropic subspaces
$Y_{k_{1}}\subseteq Y_{k_{1}+k_{2}}\subseteq\cdots\subseteq Y_{k_{1}+\cdots+k_{r}}\subseteq V_{m}.$
The stabilizer of such a flag is a parabolic subgroup $\mathrm{P}$ of $\mathrm{O}(V_{m})$ whose Levi factor $\mathrm{M}$ is
$\mathrm{M}\simeq\mathrm{GL}_{k_{1}}\times\cdots\times\mathrm{GL}_{k_{r}}\times\mathrm{O}(V_{m-k_{1}-\cdots-k_{r}}),$
where each $\mathrm{GL}_{k_{i}}$ is the group of invertible linear maps on Span $\{e_{\sum_{t=1}^{i-1}k_{t}+1},\cdots,e_{\sum_{t=1}^{i}k_{t}}\}$.
Throughout the paper, we write $\mathrm{J}_{n}$ for $\mathrm{Sp}(W_{n})$ and $\widetilde{\mathrm{J}_{n}}$ for $\widetilde{\mathrm{Sp}}(W_{n})$. We also write $\mathrm{H}_{m}^{\pm}$ for $\mathrm{SO}(V_{m}^{\pm})$.
2.1.4. Genuine representations
Let $\psi$ be a non-trivial additive character of $F$, and let $\gamma_{\psi}$ be the Weil factor associated with $\psi$ (see the appendix of [Rao93]). The character $\chi_{\psi}$ of $\widetilde{\mathrm{GL}_{k}}(F)$ is defined as follows:
$\chi_{\psi}(g,\epsilon)=\epsilon\cdot\gamma_{\psi}(\det(g))^{-1},$
where $(g,\epsilon)\in\mathrm{GL}_{k}(F)\times\{\pm 1\}$.
For a representation $\tau$ of $\mathrm{GL}_{k}(F)$, the genuine representation $\tau_{\psi}$ of $\widetilde{\mathrm{GL}_{k}}(F)$ is defined by
$\tau_{\psi}(a,\epsilon)=\epsilon\gamma_{\psi}(\det(a))\tau(a),\quad(a,\epsilon)\in\mathrm{GL}_{k}(F)\times\{\pm 1\}.$
Every irreducible admissible genuine representation of $\widetilde{\mathrm{GL}_{k}}(F)$ corresponds to a representation of $\mathrm{GL}_{k}(F)$ in this way. On the other hand, an irreducible representation $\widetilde{\pi}$ of $\widetilde{\mathrm{J}_{n}}(F)$ is called a genuine representation if $-1\in\widetilde{\mathrm{J}_{n}}(F)$ does not act trivially.
Throughout this paper, when discussing irreducible representations of $\widetilde{\mathrm{J}_{n}}(F)$ or $\widetilde{\mathrm{GL}_{k}}(F)$, we implicitly assume that they are genuine representations.
2.2. Generic characters and generic representations
Let $\mathrm{U}$ be the unipotent radical of a Borel subgroup $\mathrm{B}=\mathrm{T}\mathrm{U}$ of $\mathrm{H}_{m}^{+}$, where $\mathrm{T}$ is the $F$-rational torus stabilizing the lines $Fv_{i}$ for each $i=1,\cdots,m$. Let $\mu$ be a generic character of $\mathrm{U}(F)$ defined by
$\mu(u):=\psi(\langle ue_{2},e_{1}^{*}\rangle_{V_{m}}+\cdots+\langle ue_{m},e_{m-1}^{*}\rangle_{V_{m}}+\langle ue,e_{m}^{*}\rangle_{V_{m}}),\quad u\in\mathrm{U}(F).$
By definition, a representation $\pi$ of $\mathrm{H}_{m}^{+}$ is generic if $\operatorname{Hom}_{\mathrm{U}}(\pi,\mu)\neq 0$. Similarly, let $\mathrm{U}^{\prime}$ be the unipotent radical of the Borel subgroup $\widetilde{\mathrm{B}}^{\prime}=\widetilde{\mathrm{T}}^{\prime}\mathrm{U}^{\prime}$ of $\widetilde{\mathrm{J}_{n}}$, where $\mathrm{T}^{\prime}$ is the $F$-split torus stabilizing the lines $Fw_{i}$ for each $i=1,\cdots,n$. For each $\lambda\in F^{\times}$, define a character $\mu_{\lambda}^{\prime}$ of $\mathrm{U}^{\prime}$ by
$\mu_{\lambda}^{\prime}(u^{\prime}):=\psi(\langle u^{\prime}f_{2},f_{1}^{*}\rangle_{W_{n}}+\cdots+\langle u^{\prime}f_{n},f_{n-1}^{*}\rangle_{W_{n}}+\frac{\lambda}{2}\cdot\langle u^{\prime}f_{n}^{*},f_{n}^{*}\rangle_{W_{n}}),\quad u^{\prime}\in\mathrm{U}^{\prime}(F).$
There is an action of $\mathrm{T}(F)$ (resp. $\mathrm{T}^{\prime}(F)$) on the set of generic characters of $\mathrm{U}(F)$ (resp. $\mathrm{U}^{\prime}(F)$), where the action is given by conjugation. It is easily checked that there is a unique $\mathrm{T}(F)$-orbit of generic characters of $\mathrm{U}(F)$, and thus, the definition of a generic representation of $\mathrm{H}_{m}^{+}(F)$ does not depend on the choice of generic character $\mu$. However, in the metaplectic cases, the $\mathrm{T}^{\prime}(F)$-orbits of generic characters of $\mathrm{U}^{\prime}(F)$ are indexed by non-trivial characters of $F^{\times}$ modulo the action of $F^{\times 2}$. More precisely, the map $\lambda\rightarrow\mu_{\lambda}^{\prime}$ gives a bijection (depending on $\psi$):
$F^{\times}/F^{\times 2}\rightarrow\text{\{$\mathrm{T}^{\prime}(F)$-orbits of generic characters of $\mathrm{U}^{\prime}(F)$\}}$
(see [GGP, Sect. 12]). Therefore, if $\operatorname{Hom}_{\mathrm{U}^{\prime}}(\widetilde{\pi},\mu_{\lambda}^{\prime})\neq 0$, we say that $\widetilde{\pi}$ is $\mu_{\lambda}^{\prime}$-generic. Though the notion of genericity depends on $\psi$, when the choice of $\psi$ is clear in the context, we omit $\psi$.
We write $\operatorname{Irr}_{\text{gen}}(\widetilde{\mathrm{J}_{n}})$ for the subset of $\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$ consisting of $\mu_{1}^{\prime}$-generic representations. Often, we call an element in $\operatorname{Irr}_{\text{gen}}(\widetilde{\mathrm{J}_{n}})$ a generic representation of $\widetilde{\mathrm{J}_{n}}(F)$. The notion of generic characters and generic automorphic representations can be similarly defined for global fields.
2.3. $\gamma$-factors
Let $\mathrm{G}$ be either $\mathrm{H}_{m}=\mathrm{SO}_{2m+1}$ or $\widetilde{\mathrm{J}_{n}}=\widetilde{\mathrm{Sp}}_{2n}$ Let $\pi$ be an irreducible generic representation of $\mathrm{G}(F)$ and $\sigma$ an irreducible generic representation of $\mathrm{GL}_{l}(F)$. The local $\gamma$-factors can be defined in two distinct ways. The first approach is through the Langlands–Shahidi method ([Sha90], [L15], [Sz10]), while the second approach relies on the proportionality between the functional equations of local Rankin–Selberg type integrals (or Shimura type integrals) for $\mathrm{H}_{m}\times\mathrm{GL}_{l}$ (or $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$). It is a product of a monomial of $t=q^{-s}$ and a rational function $\frac{Q(t)}{P(t)}$, where $Q(t),P(t)\in F[t]$ such that $Q(0)=P(0)=1$. For a precise definition, please refer to [So93] and [Kap15]. It is worth noting that the two definitions of $\gamma$-factors are reconciled by several properties, as explained in [Kap15]. Since we consider both the cases of $\operatorname{char}(F)=0$ and $\operatorname{char}(F)=p$ simultaneously, we adopt the definition of $\gamma$-factors from the Langlands–Shahidi method, as they are defined in both cases. It is worthy to mention that Langlands-Shahidi method produces not only $\gamma$-factors but also local $L$-factors as follows:
Write $\gamma(s,\pi\times\sigma,\psi)=\alpha\cdot t^{k}\cdot\frac{Q(t)}{P(t)}$ for some $\alpha\in F^{\times}$ and polynomials $Q(t),P(t)\in F[t]$ such that $\text{gcd}\big{(}Q(t),P(t)\big{)}=1$ and $Q(0)=P(0)=1$. Then $L_{\psi}(s,\pi\times\sigma,\psi)$ is defined by
(2.1) $L_{\psi}(s,\pi\times\sigma)\coloneqq Q(t)^{-1}$
We would like to recall some important properties of the $\gamma$-factors that will be utilized later in our discussion. (Caution : the property (vi) does not exist in the $\operatorname{char}(F)=p$ case.)
Proposition 2.1.
Let $\pi$ and $\sigma$ be an irreducible admissible generic representation of $\mathrm{G}(F)$ and $\mathrm{GL}_{l}(F)$, respectively. Then the following statements hold:
  1. $(\mathrm{i})$
[…]
  2. $(\mathrm{ii})$
    (Multiplicity) Let $\tau_{1}\otimes\cdots\otimes\tau_{r}\otimes\pi_{0}$ (resp. $(\tau_{1})_{\psi}\otimes\cdots\otimes(\tau_{r})_{\psi}\otimes\widetilde{\pi_{0}}$) be an irreducible admissible generic representation of $\mathrm{GL}_{r_{1}}(F)\times\cdots\times\mathrm{GL}_{r_{r}}(F)\times\mathrm{H}_{k}(F)$ (resp. $\widetilde{\mathrm{GL}_{r_{1}}}\times_{\{\pm 1\}}(F)\cdots\times_{\{\pm 1\}}\widetilde{\mathrm{GL}_{r_{r}}}(F)\times_{\{\pm 1\}}\widetilde{\mathrm{J}}_{k}(F)$). Suppose that $\pi$ is the irreducible admissible generic constituent of a parabolically induced representation from $\tau_{1}\otimes\cdots\otimes\tau_{r}\otimes\pi_{0}$ or $(\tau_{1})_{\psi}\otimes\cdots\otimes(\tau_{r})_{\psi}\otimes\widetilde{\pi_{0}}$. Then,
    $\gamma(s,\pi\times\sigma,\psi)=\gamma(s,\pi_{0}\times\sigma,\psi)\prod_{i=1}^{r}\gamma(s,\tau_{i}\times\sigma,\psi)\gamma(s,\tau_{i}^{\vee}\times\sigma,\psi).$
    Here, $\gamma(s,\tau_{i}\times\sigma,\psi)$ is the $\gamma$-factor for $\mathrm{GL}_{r_{i}}\times\mathrm{GL}_{k}$ of Jacquet, Piatetski-Shapiro and Shalika [JSS].
  3. $(\mathrm{iii})$
    (Dependence on $\psi$) Given $a\in F^{\times}$, denote by $\psi_{a}$ the character of $F$ given by $\psi_{a}(x)\coloneqq\psi(ax)$ for $x\in F$. Let $\omega_{\pi},\omega_{\sigma}$ be the central character of $\pi$ and $\sigma$, respectively. Then,
    $\gamma(s,\pi\times\sigma,\psi_{a})=\omega_{\sigma}(a)^{l}|a|_{F}^{hl(s-\frac{1}{2})}\gamma(s,\pi\times\sigma,\psi).$
    Here, $h=2m$ if $\mathrm{G}=\mathrm{H}_{m}$; $h=2n$ if $\mathrm{G}=\widetilde{\mathrm{J}_{n}}$.
  4. $(\mathrm{iv})$
    (Unramified factors) When all data are unramified, we have
    $\gamma(s,\pi\times\sigma,\psi)=\frac{L_{\psi}(1-s,\pi^{\vee}\times\sigma^{\vee})}{L_{\psi}(s,\pi\times\sigma)}.$
  5. $(\mathrm{v})$
    (Global property: Functional equation) Let $K$ be a global field with a ring of adeles $\mathbb{A}$ and $\Psi$ be a nontrivial character of $K\backslash\mathbb{A}$. Assume that $\Pi$ and $\Sigma$ are generic (depending on $\Psi$) cuspidal representations of $\mathrm{G}(\mathbb{A})$ and $\mathrm{GL}_{l}(\mathbb{A})$, respectively. Let $S$ be a finite set of places of $F$ such that for $v\notin S$, all data are unramified. Then
    $L_{\Psi}^{S}(s,\Pi\times\Sigma)=\prod_{v\in S}\gamma(s,\Pi_{v}\times\Sigma_{v})\cdot L_{\Psi}^{S}(1-s,\Pi^{\vee}\times\Sigma^{\vee}).$
    Here, $L_{\Psi}^{S}(s,\Pi\times\Sigma):=\displaystyle\prod_{v\notin S}L_{\Psi_{v}}(s,\Pi_{v}\times\Sigma_{v})$ is the partial $L$-function with respect to $S$.
  6. $(\mathrm{vi})$
    (Archimedean property) For an archimedean field $F$,
    $\gamma(s,\pi\times\sigma,\psi)=\gamma^{\textrm{Artin}}(s,\pi\times\sigma,\psi).$
    Here, $\gamma^{\textrm{Artin}}(s,\pi\times\sigma,\psi)$ is the Artin $\gamma$-factor under the local Langlands correspondence.
  7. $(\mathrm{vii})$
    (Tempered $L$-function) Let $\pi,\sigma$ be irreducible tempered representations of $\mathrm{G}(F)$ and $\mathrm{GL}_{k}(F)$. Then $L_{\psi}(s,\pi\times\rho)$ is holomorphic for $\operatorname{Re}(s)>0$.
When $\mathrm{G}=\mathrm{H}_{m}\text{ and }\widetilde{\mathrm{J}_{n}}$, properties (i)-(vi) are proved in [Kap15, Theorem 1] (see also [Sha90, Theorem 3.5]) for $\operatorname{char}(F)=0$ case. However, property (vii) is only proved for $\mathrm{G}=\mathrm{H}_{m}$ ([CS, page 573], [HO13, Theorem 1.1]) in $\operatorname{char}{F}=0$ case and in [L15, Theorem 1.4] for $\operatorname{char}(F)=p$ case.
The property (vii) is proved in [CS, page 573], [HO13, Theorem 1.1] for $\operatorname{char}{F}=0$ case and in [CGL24, Theorem 1.1] for $\operatorname{char}{F}=p$ case.
Remark 2.2.
While the definition and standard properties of $\gamma$-factors for metaplectic groups in characteristic $\operatorname{char}(F)=p$ have not yet been firmly established, it is very reasonable to expect that the methods introduced in [L15] naturally extend to this case. However, due to the focus and limitations of this paper, we will adopt the working hypothesis that such extensions exist without delving into the details.
Working Hypothesis.
The $\gamma$-factors for $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$ are properly defined in $\operatorname{char}(F)=p$ cases. Furthermore, they satisfy properties (i)-(v) and (vii) in Proposition 2.1.
Remark 2.3.
In the case when $\operatorname{char}(F)=0$, Shahidi ([Sha90]) was the first to observe these properties for the Langlands–Shahidi $\gamma$-factors. Lapid–Rallis ([LR05]) and W.T. Gan ([Gan]) showed that these properties of $\gamma$-factors uniquely determine them. Therefore, when considering the Langlands–Shahidi $\gamma$-factors for $\mathrm{H}_{n}\times\mathrm{GL}_{1}$ and $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{1}$, they coincide with the $\gamma$-factors defined by Lapid–Rallis and Gan (see [Kap15, Remark 4.1]). More recently, a significant advancement has been made by Cai, Friedberg, and Kaplan ([CFK22]) in generalizing Lapid-Rallis’s doubling method to $\mathrm{G}_{n}\times\mathrm{GL}_{k}$ for split classical groups $\mathrm{G}_{n}=\mathrm{SO}(n),\mathrm{Sp}(2n)$ and $k\geq 1$. They have demonstrated that the three different definitions of $\gamma$-factors, namely the Langlands–Shahidi $\gamma$-factors, the doubling method of Lapid-Rallis and Gan, and their own generalization, actually coincide for generic representations. It is noteworthy that their generalized $\gamma$-factors are defined not only for generic representations but also for non-generic representations of $\mathrm{G}_{n}(F)\times\mathrm{GL}_{k}(F)$ in the case of $\operatorname{char}(F)=0$.
2.4. Reduction Theorem 1.1 to $\lambda=1$ cases
In this subsection, our aim is to establish a reduction from Theorem 1.1 to Theorem 2.5, which specifically focuses on $\mu_{1}^{\prime}$-generic representations of $\widetilde{\mathrm{J}_{n}}(F)$.
In order to accomplish our goal, we rely on the following lemma.
Lemma 2.4.
A $\mathrm{T}^{\prime}(F)$-orbit of irreducible $\mu_{\lambda}^{\prime}$-generic representations with respect to $\psi$ is equals to a $\mathrm{T}^{\prime}(F)$-orbit of irreducible admissible $\mu_{1}^{\prime}$-generic representations with respect to $\psi_{\lambda}$.
Proof.
Let $\pi$ be an irreducible $\mu_{\lambda}^{\prime}$-generic representation of $\widetilde{\mathrm{J}_{n}}(F)$ with respect to $\psi$. We aim to show that $\pi$ is $\mu_{1}^{\prime}$-generic with respect to $\psi_{\lambda}$.
Consider the action of $t\in\mathrm{T}^{\prime}(F)$ on a generic character $\chi^{\prime}$ of $\mathrm{U}^{\prime}(F)$, given by $(\chi^{\prime})^{t}(u^{\prime})=\chi^{\prime}(t^{-1}u^{\prime}t)$. Specifically, for $t\in\mathrm{T}^{\prime}(F)$ with $t(f_{i})=t_{i}f_{i}$ and $t(f_{i}^{*})=t_{i}^{-1}f_{i}^{*}$ $(1\leq i\leq n)$, we have
$(\mu_{\lambda}^{\prime})^{t}(u^{\prime})=\psi\left(\sum_{i=1}^{m-1}t_{i}^{-1}t_{i+1}\langle u^{\prime}f_{i},f_{i+1}^{*}\rangle+\lambda t_{m}^{-2}\frac{\langle u^{\prime}f_{m}^{*},f_{m}^{*}\rangle}{2}\right).$
In particular, if we choose $\underline{t}\in\mathrm{T}^{\prime}(F)$ specifically with $t_{i}=\lambda^{-(m-i)}$, then we obtain
$(\mu_{\lambda}^{\prime})^{\underline{t}}(u^{\prime})=\psi_{\lambda}\left(\sum_{i=1}^{m-1}\langle u^{\prime}f_{i},f_{i+1}^{*}\rangle+\frac{\langle u^{\prime}f_{m}^{*},f_{m}^{*}\rangle}{2}\right).$
Therefore, a $(\mu_{\lambda}^{\prime})^{\underline{t}}$-generic representation (with respect to $\psi$) is $\mu_{1}^{\prime}$-generic with respect to $\psi_{\lambda}$. Since the notion of genericity is invariant under $\mathrm{T}^{\prime}(F)$-orbits, it follows that $\pi$ is $\mu_{1}^{\prime}$-generic with respect to $\psi_{\lambda}$.
This completes the proof. ∎
The lemma above and Proposition 2.1 (iii) imply that it suffices to consider $\lambda=1$ cases to prove Theorem 1.1. Hence, for the remainder of the paper, we will focus on establishing the following theorem, which serves as a replacement for Theorem 1.1.
Theorem 2.5 (Local Converse Theorem for $\widetilde{\mathrm{Sp}}_{2n}$).
Let $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ be irreducible admissible $\mu_{1}^{\prime}$-generic representations of $\widetilde{\mathrm{J}_{n}}(F)$ such that
$\gamma(s,\widetilde{\pi}\times\sigma,\psi)=\gamma(s,\widetilde{\pi}^{\prime}\times\sigma,\psi)$
holds for all irreducible generic supercuspidal representation $\sigma$ of $\mathrm{GL}_{i}(F)$ with $1\leq i\leq n$. Then
$\widetilde{\pi}\simeq\widetilde{\pi}^{\prime}.$
3. Local theta correspondence
In this section, we introduce the local theta correspondence induced by a Weil representation $\omega_{\psi,V_{m},W_{n}}$ of $\mathrm{H}_{m}(F)\times\widetilde{\mathrm{J}}_{n}(F)$ and recapitulate some fundamental results associated with it. Subsequently, we prove two theorems that will play crucial roles in establishing our main theorem in the next section.
The group $\mathrm{H}_{m}(F)\times\widetilde{\mathrm{J}}_{n}(F)$ naturally possesses a distinguished representation known as the Weil representation, denoted as $\omega=\omega_{\psi,V_{m},W_{n}}$. For every $\pi\in\operatorname{Irr}(\mathrm{H}_{m})$, the maximal $\pi$-isotypic quotient of $\omega$ can be expressed as
$\pi\boxtimes\Theta_{\psi,W_{n},V_{m}}(\pi),$
where $\Theta_{\psi,W_{n},V_{m}}(\pi)$ denotes a smooth finite-length representation of $\widetilde{\mathrm{J}_{n}}(F)$ referred to as the big theta lift of $\pi$. The small theta lift of $\pi$ corresponds to the maximal semisimple quotient of $\Theta_{\psi,W_{n},V_{m}}(\pi)$. Similarly, for each $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$, we obtain a smooth finite-length representation $\Theta_{\psi,V_{m},W_{n}}(\widetilde{\pi})$ of $\mathrm{H}_{m}(F)$. The maximal semisimple quotient of $\Theta_{\psi,V_{m},W_{n}}(\widetilde{\pi})$ (and $\Theta_{\psi,W_{n},V_{m}}(\pi)$) is denoted as $\theta_{\psi,V_{m},W_{n}}(\widetilde{\pi})$ (and $\theta_{\psi,W_{n},V_{m}}(\pi)$) respectively. The Howe duality conjecture, originally proven by Waldspurger [Wa90] for non-residue characteristic $2$ and later by Gan–Takeda [GT1, GT2] in a more general setting, asserts that if these small theta lifts are nonzero, then they are irreducible.
For the rest of this paper, we focus exclusively on the case where $m=n$ and conveniently denote $V_{n}$ and $W_{n}$ as $V$ and $W$ respectively. Given $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$, there exists a unique $\epsilon\in\{\pm 1\}$ such that $\theta_{\psi,V^{\epsilon},W}(\widetilde{\pi})$ is nonzero (see [GS, Theorem 1.4]).
Using the theta correspondence between $\widetilde{\mathrm{J}_{n}}$ and $\mathrm{H}_{n}$, we define a mapping $\Theta_{\psi}:\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})\longrightarrow\operatorname{Irr}(\mathrm{H}_{n}^{+})\sqcup\operatorname{Irr}(\mathrm{H}_{n}^{-})$ for each non-trivial additive character $\psi:F\longrightarrow\mathbb{C}^{\times}$. Specifically, for $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$, we have $\Theta_{\psi}(\widetilde{\pi})=\theta_{\psi,V^{\epsilon},W}(\widetilde{\pi})$, where $\epsilon\in\{\pm 1\}$ is chosen uniquely such that $\theta_{\psi,V^{\epsilon},W}(\widetilde{\pi})$ is nonzero. The inverse of this map is denoted by $\Theta_{\psi}^{-1}(\pi)=\theta_{\psi,W,V}(\pi)$, which is nonzero for any $\pi\in\text{Irr}(\mathrm{H}_{n}^{\pm})$.
The bijection $\Theta_{\psi}$ preserves various properties. Here are some of them that will be useful later.
Theorem 3.1 ([GI1, Theorem C.4]).
Let $\pi\in\operatorname{Irr}_{\operatorname{temp}}(\mathrm{H}_{n})$ and $\widetilde{\pi}\in\operatorname{Irr}_{\operatorname{temp}}(\widetilde{\mathrm{J}_{n}})$. Then $\Theta_{\psi,W,V}(\pi)\in\operatorname{Irr}_{\operatorname{temp}}(\widetilde{\mathrm{J}_{n}})$. Similarly, if $\Theta_{\psi,V,W}(\widetilde{\pi})$ is non-zero, then $\Theta_{\psi,V,W}(\widetilde{\pi})\in\operatorname{Irr}_{\operatorname{temp}}(\mathrm{H}_{n})$.
Theorem 3.2 ([JS03, Theorem 2.2]).
Let $\pi\in\operatorname{Irr}_{\operatorname{supc}}(\mathrm{H}_{n})$ and $\widetilde{\pi}\in\operatorname{Irr}_{\operatorname{supc}}(\widetilde{\mathrm{J}_{n}})$. Suppose that $\pi$ is generic and $\widetilde{\pi}$ is $\mu_{1}^{\prime}$-generic. Then $\Theta_{\psi,W,V}(\pi)$ is supercuspidal and
$\Theta_{\psi,V,W}(\widetilde{\pi})=\begin{cases}\mathrm{Steinberg\ representation},\ \mathrm{if\ }n=1\mathrm{\ and\ }\widetilde{\pi}\mathrm{\ is\ odd\ Weil\ representation\ }\\ \mathrm{supercuspidal},\quad\quad\quad\quad\ \ \mathrm{\ otherwise.}\end{cases}$
The following theorem extends [JS03, Corollary 2.1] from supercuspidal representations to tempered representations.
Theorem 3.3.
Let $\widetilde{\pi}\in\operatorname{Irr}_{\operatorname{temp}}(\widetilde{\mathrm{J}_{n}})$. If $\widetilde{\pi}$ is $\mu_{1}^{\prime}$-generic, then $\Theta_{\psi}(\widetilde{\pi})$ is a tempered generic representation of the split odd orthogonal group $\mathrm{H}_{n}^{+}(F)$.
Indeed, this is a part of [GS, Theorem 9.1]. However, unfortunately, the proof is not provided there. To show that their omitted proof also works for the case $\operatorname{char}(F)=p$, we provide a proof here. The proof relies on the following lemma, which computes the twisted Jacquet module of the Weil representation, as stated in [GS, Proposition 9.2] without proof.
Lemma 3.4 ([GS, Proposition 9.2]).
Let $(\omega_{\psi,V,W})_{\mathrm{U},\mu^{-1}}$ be the twisted Jacquet module of $\omega_{\psi,V,W}$ with respect to $\mathrm{U}$ and $\mu^{-1}$. Then the following holds:
$(\omega_{\psi,V,W})_{\mathrm{U},\mu^{-1}}\cong\operatorname{ind}_{\mathrm{U}^{\prime}}^{\widetilde{\mathrm{J}_{n}}}(\mu_{1}^{\prime}).$
The proof of Lemma 3.4 will be provided in the next section. We can establish Theorem 3.3 assuming Lemma 3.4 as follows.
Proof of Theorem 3.3.
Write $(\omega_{\psi,V,W})_{\mathrm{U},\mu^{-1}}$ for the twisted Jacquet module of $\omega_{\psi,V,W}$ with respect to $\mathrm{U}$ and $\mu^{-1}$. By Lemma 3.4, we have
$\operatorname{Hom}_{\widetilde{\mathrm{J}_{n}}\times\mathrm{U}}\big{(}\omega_{\psi,V,W},\ \widetilde{\pi}\otimes\mu^{-1}\big{)}$ $\cong\operatorname{Hom}_{\widetilde{\mathrm{J}_{n}}}\big{(}(\omega_{\psi,V,W})_{\mathrm{U},\mu^{-1}},\ \widetilde{\pi}\big{)}$
$\cong\operatorname{Hom}_{\widetilde{\mathrm{J}_{n}}}\big{(}\operatorname{ind}_{\mathrm{U}^{\prime}}^{\widetilde{\mathrm{J}_{n}}}(\mu_{1}^{\prime}),\ \widetilde{\pi}\big{)}$
$\cong\operatorname{Hom}_{\mathrm{U}^{\prime}}\big{(}\mu_{1}^{\prime},\ (\widetilde{\pi}^{\vee}\mid_{\mathrm{U}^{\prime}})^{\vee}\big{)}$
$\cong\operatorname{Hom}_{\mathrm{U}^{\prime}}\big{(}\widetilde{\pi}^{\vee},\ (\mu_{1}^{\prime})^{\vee}\big{)}$
$\cong\operatorname{Hom}_{\mathrm{U}^{\prime}}\big{(}\widetilde{\pi},\ \mu_{1}^{\prime}\big{)},$
where the last equality follows from the facts that $\widetilde{\pi}$ and $\mu_{1}^{\prime}$ are unitary.
On the other hand,
$\operatorname{Hom}_{\widetilde{\mathrm{J}_{n}}\times\mathrm{U}}\big{(}\omega_{\psi,V,W},\ \widetilde{\pi}\otimes\mu^{-1}\big{)}\cong\operatorname{Hom}_{\mathrm{U}}\big{(}\Theta_{\psi,V,W}(\widetilde{\pi}),\ \mu^{-1}\big{)}.$
If $\widetilde{\pi}^{\prime}$ is $\mu_{1}^{\prime}$-generic, then $\operatorname{Hom}_{\mathrm{U}^{\prime}}\big{(}\widetilde{\pi}^{\prime},\ \mu_{1}^{\prime}\big{)}\neq 0$, and thus $\operatorname{Hom}_{\mathrm{U}}\big{(}\Theta_{\psi,V,W}(\widetilde{\pi}^{\prime}),\ \mu^{-1}\big{)}\neq 0$. This, combined with Theorem 3.1, proves the theorem and $\Theta_{\psi,V,W}(\widetilde{\pi}^{\prime})$ is a representation of $\mathrm{H}_{n}^{+}$ because it is generic. ∎
Remark 3.5.
The opposite direction of Theorem 3.3, that is starting with a tempered generic representation of $\mathrm{H}_{n}(F)$ and its theta lift to $\widetilde{\mathrm{J}_{n}}(F)$ is generic, is proved in greater generality in [JS03, Corollary 2.1 (1)]. Note that when $n=1$, even though $\widetilde{\pi}$ is $\mu_{1}^{\prime}$-generic and a supercuspidal representation, $\Theta_{\psi}(\widetilde{\pi}^{\prime})$ might not be supercuspidal. For example, if $\widetilde{\pi}$ is the odd Weil representation, then $\Theta_{\psi}(\widetilde{\pi}^{\prime})$ is a Steinberg representation. For this reason, Jiang and Soudry separated the proof of the LCT for $\mathrm{H}_{n}$ according to whether $n=1$ or $n\geq 2$. In this paper, to ensure our arguments also work for $n=1$ cases, we reduce the proof of our main theorem to the tempered cases, not the supercupidal cases. This contrasts with the approach taken in [JS03], where they reduced it to the supercuspidal cases. This is one of the delicate parts in this paper.
The following is a special case of Kudla’s supercuspidal support theorem.
Theorem 3.6 ([Ku, Theorem 7.1] or [GI1, Proposition 5.2]).
Let $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$ be a $\mu_{1}^{\prime}$-generic representation. Suppose that $\Theta_{\psi,V,W}(\widetilde{\pi})$ is nonzero. Put $\pi=\Theta_{\psi,V,W}(\widetilde{\pi})$. Assume that $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$ and $\pi\in\operatorname{Irr}(\mathrm{H}_{n})$ have supercuspidal supports $(\widetilde{\rho_{1}},\cdots,\widetilde{\rho_{r}};\widetilde{\pi_{0}})$ with $\widetilde{\pi_{0}}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n_{0}}})$ and $(\rho_{1},\cdots,\rho_{s};\pi_{0})$ with $\pi_{0}\in\operatorname{Irr}(\mathrm{H}_{m_{0}})$, respectively. Then the following holds:
  1. $(\mathrm{i})$
    $m_{0}=n_{0}$ or $m_{0}=n_{0}-1$.
  2. $(\mathrm{ii})$
    $\pi_{0}=\theta_{\psi,V_{m_{0}},W_{n_{0}}}(\widetilde{\pi_{0}})$.
  3. $(\mathrm{iii})$
    $\{\rho_{1},\cdots,\rho_{s}\}=\begin{cases}\{\widetilde{\rho_{1}},\cdots,\widetilde{\rho_{r}}\},\quad\quad\quad\ \mathrm{if}\ m_{0}=n_{0}\\ \{\widetilde{\rho_{1}},\cdots,\widetilde{\rho_{r}},|\cdot|^{\frac{1}{2}}\},\quad\mathrm{if}\ m_{0}=n_{0}-1\end{cases}$
By applying Kudla’s supercuspidal theorem, we can establish that $\Theta_{\psi}$ preserves $\gamma$-factors.
Theorem 3.7.
For $\widetilde{\pi}\in\operatorname{Irr}_{\operatorname{gen}}(\widetilde{\mathrm{J}_{n}})$, suppose that $\pi=\Theta_{\psi}(\widetilde{\pi})$ is generic. Then for any irreducible generic representation $\sigma$ of $\mathrm{GL}_{r}(F)$,
$\gamma(s,\widetilde{\pi}\times\sigma,\psi)=\gamma(s,\pi\times\sigma,\psi).$
Remark 3.8.
This result may follow from [GS, Proposition 11.1] and the discussion after [GS, Corollary 11.2]. In the proof provided by [GS, Proposition 11.1], they construct first $\Pi$ (the globalization of $\pi$) and consider its first occurrence, denoted as $\Theta_{k_{0}}(\Pi)$, in the theta lifting tower from $\mathrm{H}_{n}(\mathbb{A})$ to the tower of $\widetilde{\mathrm{J}_{2k}}(\mathbb{A})$. They then compare the $\gamma$-factors of $\Pi_{v}$ and $\Theta_{k_{0}}(\Pi)_{v}$ at unramified places. However, the genericity of $\Theta_{k_{0}}(\Pi)_{v}$ remains uncertain, as there is no guarantee that $\Theta_{k_{0}}(\Pi)$ is globally generic. Since local $\gamma$-factors of $\widetilde{\mathrm{J}_{k_{0}}}(F_{v})$ are only defined for generic representations, we cannot directly compare the local $\gamma$-factors of $\Pi_{v}$ and $\Theta_{k_{0}}(\Pi)_{v}$ if $\Theta_{k_{0}}(\Pi)_{v}$ happens to be non-generic. Therefore, we provide a detailed proof circumventing this issue.
Proof.
When $r=1$, the theorem is incorporated into [GI1, Theorem 11.5] in $\operatorname{char}(F)=0$ case. In the case of $\operatorname{char}(F)=p\neq 2$, a parallel reasoning to the proof of [GI1, Theorem 11.5] can be applied, utilizing the global-to-local argument from [GI1, subsection 11.6]. Notably, in this scenario, [GL, Theorem 1.1] plays a comparable role. Consequently, we may presume the validity of [GI1, Theorem 11.5] across all characteristics except 2.
However, when considering the general case of $r\neq 1$, there is a significant difference. While [GI1, Theorem 11.5] is formulated using the Lapid-Rallis and W.T. Gan’s $\gamma$-factors, which are defined for all irreducible admissible representations of $\widetilde{\mathrm{J}_{n}}(F)\times\mathrm{GL}_{1}(F)$ and $\mathrm{H}_{n}(F)\times\mathrm{GL}_{1}(F)$, including non-generic ones, the $\gamma$-factors appearing in the theorem are the Langlands-Shahidi’s $\gamma$-factors, which are defined only for generic representations. This distinction requires careful consideration when employing a global-to-local argument as presented below.
We begin by proving the case when $\sigma$ has a nonzero Iwahori-fixed vector. In this case, we can express $\sigma$ as a submodule of a principal series representation of $\mathrm{GL}_{r}(F)$. By utilizing the multiplicative property of $\gamma$-factors, it suffices to consider the case when $\sigma$ is a $1$-dimensional character of $F^{\times}$. Since our $\gamma$-factors coincide with Lapid-Rallis and Gan’s $\gamma$-factors (as stated in Remark 2.3), the result for this case follows from [GI1, Theorem 11.5].
Next, we consider the general case. Again, using the multiplicative property of $\gamma$-factors and Theorem 3.6, we only need to consider the case when both $\widetilde{\pi}$ and $\sigma$ are supercuspidal. At this point, we employ a global-to-local argument. Let us introduce the following data:
  • $K$: a totally imaginary number field with $\mathbb{A}$ as its adele ring such that $K_{v_{0}}=F$ for a specific finite place $v_{0}$ of $K$.
  • $\Psi$: a nontrivial additive character of $K\backslash\mathbb{A}$ satisfying $\Psi_{v_{0}}=\psi$.
  • $\mathfrak{U}^{\prime}$: a nontrivial additive generic character of $\mathrm{U}^{\prime}(K)\backslash\mathrm{U}^{\prime}(\mathbb{A})$ associated with $\Psi$, such that $\mathfrak{U}_{v_{0}}^{\prime}=\mu_{1}^{\prime}$.
  • $\mathbb{\mathbb{W}}$: a symplectic space over $K$ of dimension $2n$, with the associated isometry group $\mathrm{Sp}(\mathbb{W})$.
  • $\mathbb{V}$: a symmetric space over $K$ of dimension $12n+$ 1, such that $\mathbb{V}_{v_{0}}=V$, with the associated isometry group $\mathrm{SO}(\mathbb{V})$.
  • $\{\mathbb{V}_{k}\}_{k\geq 0}$: The Witt tower over $K$ containing $\mathbb{V}_{n}=\mathbb{V}$.
  • $\Sigma$: a globally generic cuspidal representation of $\mathrm{GL}_{r}(\mathbb{A})$, such that $\Sigma_{v_{0}}=\sigma$, and for all finite places $v\neq v_{0}$ of $K$, the local components $\Sigma_{v}$ are unramified.
The construction of such $\Sigma$ is indeed possible, as stated in [PS, Theorem 4.1] for the case when $\operatorname{char}(F)=0$ and in [GL, Theorem 1.1] for the case when $\operatorname{char}(F)=p$.
In order to prove the desired result for the case when $\pi$ is supercuspidal, we construct a globally $\mathfrak{U}^{\prime}(\mathbb{A})$-generic cuspidal representation $\Pi$ of $\widetilde{\mathrm{Sp}}(\mathbb{W})(\mathbb{A})$, adapting the constructions from [PS, Theorem 4.1] and [GL, Theorem 1.1] to metaplectic groups. This representation $\Pi$ satisfies $\Pi_{v_{0}}=\widetilde{\pi}$ and for all finite places $v\neq v_{0}$ of $K$, $\Pi_{v}$ is unramified.
Let $\Xi$ be the global theta lift of $\Pi$ to $\mathrm{SO}(\mathbb{V})(\mathbb{A})$. Using [Fu, Proposition 3], we know that $\Xi$ is nonzero and generic. We claim that $\Xi$ is cuspidal. Suppose not, which means that there exists an integer $m<n$ and a $(2m+1)$-dimensional symmetric space $\mathbb{V}_{m}$ over $K$ such that the global theta lift $\Theta_{m}(\Pi)$ of $\Pi$ to $\mathrm{SO}(\mathbb{V}_{m})(\mathbb{A})$ is cuspidal. Let $\Theta_{m}(\Pi)_{0}$ be an irreducible summand of $\Theta_{m}(\Pi)$. Then $\Theta_{m}(\Pi)_{0,v_{0}}$ is the nonzero local theta lift of $\widetilde{\pi}$ to $\mathrm{SO}(V_{m})$. However, this contradicts the tower property of local theta lifts (see [Ku, Theorem 6.1]) since $\widetilde{\pi}$ is a supercuspidal representation of $\mathrm{H}_{n}(F)$. Therefore, we have shown that $\Xi$ is a cuspidal representation of $\mathrm{SO}(\mathbb{V})(\mathbb{A})$ with $\Xi_{v_{0}}=\theta_{\psi,V,W}(\widetilde{\pi})$.
Let $\Xi_{0}$ be an irreducible generic constituent of $\Xi$. (Indeed, $\Xi$ is irreducible by the multiplicity one property in the Howe duality theorem.) By our first claim, for all unramified places $v$ of $K$, we have
$\gamma(s,\Xi_{0,v}\times\Sigma_{v},\Psi_{v})=\gamma(s,\Pi_{v}\times\Sigma_{v},\Psi_{v}).$
For any archimedean place $v$ of $K$, $K_{v}=\mathbb{C}$. The theta correspondence for complex groups is well-understood and can be described in terms of the local Langlands correspondence (see [AB95]). Applying the archimedean property of $\gamma$-factors, we have the above identity at the archimedean places of $K$ as well.
Put $S_{0}$ the set of all archimedean places of $K$ and $S=S_{0}\cup\{v_{0}\}$ (In $\operatorname{char}(F)=p$ cases, put $S=\{v_{0}\}$). Note that for all places $v\notin S$, $\Xi_{0,v},\Pi_{v}$ and $\Sigma_{v}$ are unramified, and for all archimedean places $v$, $K_{v}=\mathbb{C}$. By the global functional equation of $\gamma$-factors,
$1=$ $\,\prod_{v\in S}\gamma(s,\Xi_{0,v}\times\Sigma_{v},\Psi_{v})\cdot\frac{L^{S}(1-s,\Xi_{0}^{\vee}\times\Sigma^{\vee})}{L^{S}(s,\Xi_{0}\times\Sigma)}$
$=$ $\,\prod_{v\in S}\gamma(s,\Pi_{v}\times\Sigma_{v},\Psi_{v})\cdot\frac{L^{S}(1-s,\Pi^{\vee}\times\Sigma^{\vee})}{L^{S}(s,\Pi\times\Sigma)}.$
By canceling the identities at places outside $v_{0}$, we have
$\gamma(s,\Xi_{0,v_{0}}\times\Sigma_{v_{0}},\Psi_{v_{0}})=\gamma(s,\Pi_{v_{0}}\times\Sigma_{v_{0}},\Psi_{v_{0}}).$
Since $\Pi_{v_{0}}=\widetilde{\pi}^{\prime}$, $\Sigma_{v_{0}}=\sigma$, and $\Xi_{0,v_{0}}=\pi$, we get the desired identity.
Now, let’s consider the case when $\pi$ is the Steinberg representation. We choose another finite place $v_{1}\neq v_{0}$ of $K$ and a supercuspidal representation $\pi_{1}$ of $\mathrm{H}_{n}(F)$. Using similar constructions as in [PS, Theorem 4.1] and [GL, Theorem 1.1], we can construct a $\mathfrak{U}^{\prime}(\mathbb{A})$-generic cuspidal representation $\Pi^{\prime}$ of $\widetilde{\mathrm{Sp}}_{n}(\mathbb{A})$ such that $\Pi^{\prime}_{v_{0}}=\widetilde{\pi}$, $\Pi^{\prime}_{v_{1}}=\theta_{\psi,W,V}(\pi_{1})$, and it is unramified at other finite places $v\neq v_{0},v_{1}$ of $K$.
Let $\Xi_{0}$ be an irreducible generic constituent of the global theta lift $\Theta(\Pi^{\prime})$ of $\Pi^{\prime}$ to $\mathrm{SO}(\mathbb{V})(\mathbb{A})$. Then, $\Xi_{0,v_{0}}\simeq\pi$, $\Xi_{0,v_{1}}\simeq\pi_{1}$, and so $\Xi_{0}$ is cuspidal because $\pi_{1}$ is supercuspidal.
We have already shown that
$\gamma(s,\Xi_{0,v_{1}}\times\Sigma_{v_{1}},\Psi_{v_{1}})=\gamma(s,\Pi_{v_{1}}^{\prime}\times\Sigma_{v_{1}},\Psi_{v_{1}})$
since $\Xi_{0,v_{1}}\simeq\pi_{1}$ is supercuspidal.
For other places $v\neq v_{0},v_{1}$, we know that
$\gamma(s,\Xi_{0,v}\times\Sigma_{v},\Psi_{v})=\gamma(s,\Pi_{v}^{\prime}\times\Sigma_{v},\Psi_{v})$
from the previous argument.
Therefore, by canceling the above identities at all places other than $v=v_{0}$ in the two global functional equations, we obtain the desired identity. ∎
4. Computation of the twisted Jacquet module of the Weil representation
In this section, we establish Lemma 3.4, which plays a crucial role in proving Theorem 3.3.
Let $V=V_{n}$ and $W=W_{n}$, and consider a symplectic form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) defined on $V\otimes W$ as follows:
$(v_{1}\otimes w_{1},v_{2}\otimes w_{2})=\langle v_{1},v_{2}\rangle_{V}\cdot\langle w_{1},w_{2}\rangle_{W}.$
There is a natural embedding of $\mathrm{H}_{n}\times\mathrm{J}_{n}$ (resp. $\mathrm{H}_{n}\times\widetilde{\mathrm{J}_{n}}$) into $\mathrm{Sp}(V\otimes W)$. By restricting the Weil representation $\omega_{\psi,V,W}$ of $\widetilde{\mathrm{Sp}}(V\otimes W)(F)$ to the image of $\mathrm{H}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$ inside $\widetilde{\mathrm{Sp}}(V\otimes W)(F)$, we regard $\omega_{\psi,V,W}$ as a representation of $\mathrm{H}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$.
Recall that $V=Y_{n}\oplus e\oplus Y_{n}^{*}$ and $W=X_{n}\oplus X_{n}^{*}$. There are two kinds of polarization for $V\otimes W$. One is $(Y_{n}^{*}\otimes W)\oplus(e\otimes X_{n})$, and the other is $V\otimes X_{n}^{*}$. Depending on the chosen polarization, the description of the associated Weil representation will differ. For our purpose, we use the polarization $(Y_{n}^{*}\otimes W)\oplus(e\otimes X_{n})$ of $V\otimes W$. In this case, there is an action of $\mathrm{H}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$ (resp. $\widetilde{\mathrm{J}_{n}}(F)$) on the Schwartz-Bruhat function space $\big{(}S(Y_{n}^{*}\otimes W)\otimes S(e\otimes X_{n})\big{)}(F)$ (resp. $S(X_{n})(F)$) via the mixed Schrödinger model of the Weil representation $\omega_{\psi,V,W}$ (resp. $\omega_{\psi,W}$).
Let $\mathrm{P}_{n}=\mathrm{M}_{n}\mathrm{V}_{n}$ be a parabolic subgroup of $\mathrm{H}_{n}$ that stabilizes $Y_{n}$, where $\mathrm{M}_{n}$ is the Levi subgroup and $\mathrm{V}_{n}$ is the maximal unipotent subgroup of $\mathrm{M}_{n}$. Write $\mathrm{N}_{n}=\{\alpha\in\operatorname{Hom}(Y_{n}^{*},Y_{n})\ |\ \alpha^{*}=-\alpha\},$ where $\alpha^{*}$ is the element in $\operatorname{Hom}(Y_{n}^{*},Y_{n})$ satisfying
$\langle ay_{1},y_{2}\rangle_{V}=\langle y_{1},a^{*}y_{2}\rangle_{V},\quad\text{ for all }y_{1},y_{2}\in Y_{n}^{*}.$
Then
$\mathrm{M}_{n}\simeq\mathrm{GL}(Y_{n})\text{ and }\mathrm{V}_{n}\simeq\operatorname{Hom}(e,Y_{n})\ltimes\mathrm{N}_{n}.$
Let $\mathrm{Z}_{n}$ be the maximal unipotent subgroup of $\mathrm{M}_{n}$. The action of $\mathrm{Z}_{n}(F)\cdot\operatorname{Hom}(e,Y_{n})\cdot\mathrm{N}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$ on $S(Y_{n}^{*}\otimes W)(F)\otimes S(e\otimes X_{n})(F)$ via the mixed Schrödinger model can be described as follows (see [MVW, §II.7, Chapter 2]). For $\phi_{1}\otimes\phi_{2}\in S(Y_{n}^{*}\otimes W)(F)\otimes S(e\otimes X_{n})(F)$ and $(w,x)\in(Y_{n}^{*}\otimes W)(F)\times(e\otimes X_{n})(F)$,
  • $\omega_{\psi,V,W}(1,h)(\phi_{1}\otimes\phi_{2})(w,x)=\phi_{1}(w\bar{h})\cdot(\omega_{\psi,W}(h)\phi_{2})(x)\text{ for }h\in\widetilde{\mathrm{J}_{n}}(F),$
  • $\omega_{\psi,V,W}\big{(}z,1\big{)}(\phi_{1}\otimes\phi_{2})(w,x)=\phi_{1}\big{(}z^{*}w\big{)}\cdot\phi_{2}(x)\text{ for }z\in\mathrm{Z}_{n}(F)\subset\mathrm{GL}(Y_{n})(F),$
  • $\omega_{\psi,V,W}(t,1)(\phi_{1}\otimes\phi_{2})(w,x)=\psi\big{(}(t^{*}w,x)\big{)}\cdot(\phi_{1}\otimes\phi_{2})(w,x)\text{ for }t\in\operatorname{Hom}(e,Y_{n})(F),$
  • $\omega_{\psi,V,W}(s,1)\cdot(\phi_{1}\otimes\phi_{2})(w,x)=\psi(\frac{1}{2}\langle sw,w\rangle_{W})\cdot(\phi_{1}\otimes\phi_{2})(w,x)\text{ for }s\in\mathrm{N}_{n}(F)\subset\operatorname{Hom}(Y_{n}^{*},Y_{n})(F),$
where $\bar{h}\text{ is the projection of }h\text{ to }\mathrm{J}_{n}(F)$ and $z^{*}\in\mathrm{GL}(Y_{n}^{*})(F),\ t^{*}\in\operatorname{Hom}(Y_{n}^{*},e)(F)$ are adjoint maps of $z$ and $t$, respectively.
For convenience, we identify $Y_{n}^{*}\otimes W$ with $W^{n}$ using the basis $\{e_{n}^{*},\cdots,e_{1}^{*}\}$ of $Y_{n}^{*}$ and $e\otimes X_{n}$ with $X_{n}$.
Let $\mathrm{M}_{n\times n}$ be the $(n\times n)$ matrix group. Using the basis $\{e_{1},\cdots,e_{n}\}$ of $Y_{n}$, $\{e_{n}^{*},\cdots,e_{1}^{*}\}$ of $Y_{n}^{*}$ and $\{f_{1},\cdots,f_{n}\}$ of $X_{n}$, we can identify $\mathrm{Z}_{n}(F)$, $\operatorname{Hom}(e,Y_{n})$ and $\mathrm{N}_{n}(F)$ as a subgroup of $\mathrm{GL}_{n}(F)$, $F^{n}$ and $\mathrm{M}_{n\times n}(F)$, respectively.
Let $\varpi$ be the $(n\times n)$ anti-diagonal matrix whose non-zero elements are all one. With the above identifications, we can describe the action of $\mathrm{Z}_{n}(F)\cdot\operatorname{Hom}(e,X_{n})\cdot\mathrm{N}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$ on $\big{(}S(W^{n})\otimes S(X_{n})\big{)}(F)$ as follows.
For $\phi_{1}\otimes\phi_{2}\in\big{(}S(W^{n})\otimes S(X_{n})\big{)}(F)$ and $(w_{1},\cdots,w_{n};x)\in(W^{n}\oplus X_{n})(F)$,
(4.1) $\omega_{\psi,V,W}(1,h)(\phi_{1}\otimes\phi_{2})(w_{1},\cdots,w_{n};x)=\phi_{1}(w_{1}\bar{h},\cdots,w_{n}\bar{h})\cdot(\omega_{\psi,W}(h)\phi_{2})(x)\text{ for }h\in\widetilde{\mathrm{J}_{n}}(F),$
(4.2) $\omega_{\psi,V,W}\big{(}z,1\big{)}(\phi_{1}\otimes\phi_{2})(w_{1},\cdots,w_{n};x)=\phi_{1}\big{(}(w_{1},\cdots,w_{n})\cdot\varpi z\varpi\big{)}\cdot\phi_{2}(x)\text{ for }z\in\mathrm{Z}_{n}(F),$
(4.3) $\omega_{\psi,V,W}(t,1)(\phi_{1}\otimes\phi_{2})(w_{1},\cdots,w_{n};x)=\psi\big{(}\sum_{i=1}^{n}t_{i}\cdot\langle w_{n+1-i},x\rangle_{W}\big{)}(\phi_{1}\otimes\phi_{2})(w_{1},\cdots,w_{n};x)\text{ for }t\in F^{n},$
(4.4) $\omega_{\psi,V,W}(s,1)(\phi_{1}\otimes\phi_{2})(w_{1},\cdots,w_{n};x)=\psi(\frac{1}{2}\mathrm{tr}\big{(}Gr(\mathbf{w})\cdot\varpi\cdot s\big{)})(\phi_{1}\otimes\phi_{2})(w_{1},\cdots,w_{n};x)\text{ for }s\in\mathrm{N}_{n}(F),$
where $\mathbf{w}=(w_{1},v_{2},\ldots,w_{n}),\ Gr(\mathbf{w})=\big{(}\langle w_{i},w_{j}\rangle_{W}\big{)}.$
Inspired by the proof of [JS03, Proposition 2.1, Corollary 2.1], we now prove Lemma 3.4. It is worth mentioning that [GS, Proposition 9.2] provides partial coverage of this theorem for $\mathrm{char}(F)=0$ without presenting a proof.
Proof of Lemma 3.4.
We choose the polarization $W^{n}\oplus X_{n}$ of $V\otimes W$ and utilize the action of the Weil representation with this polarization. Let us define
$\overline{W_{0}}=\left\{\mathbf{w}=(w_{1},\cdots,w_{n})\in W^{n}\ |\ Gr(\mathbf{w})=\begin{pmatrix}0&\cdots&0\\ \vdots&\vdots&\vdots\\ 0&\cdots&0\end{pmatrix}\right\},$
where $Gr(\mathbf{w})=\big{(}(w_{i},w_{j})\big{)}$. We first claim that
$(\omega_{\psi,V_{n},W_{n}})_{\mathrm{U},\mu^{-1}}\simeq S(\overline{W_{0}}\oplus X_{n}).$
Note that $\overline{W_{0}}$ is a closed subset of $W^{n}$. Therefore, by [BZ], we have the exact sequence
[…]
where $\overline{\text{i}}$ is induced from the open inclusion map $i:W^{n}\oplus X_{n}\backslash\overline{W_{0}}\oplus X_{n}\to W^{n}\oplus X_{n}$ and $\overline{\text{res}}:\mathcal{S}(W^{n}\oplus X_{n})\to\mathcal{S}(\overline{W_{0}}\oplus X_{n})$ is the restriction map. Let $J_{\mathrm{U},\mu^{-1}}$ be the twisted Jacquet functor with respect to $\mathrm{U}$ and $\mu^{-1}$. Since the functor $J_{\mathrm{U},\mu^{-1}}$ is exact, we have the exact sequence
[…]
Therefore, our first claim is proved. Note that there is an action of $\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$ on $\overline{W_{0}}$ inherited from $W^{n}$ as follows;
$(w_{1},\cdots,w_{n})\cdot(z,h)=((w_{1}\bar{h},\cdots,w_{n}\bar{h})\varpi z\varpi).$
Using the $\mathrm{Z}_{n}(F)$-action on $\overline{W_{0}}$, we can choose the representatives of the orbits of $\overline{W_{0}}$ as the form
$(0,\cdots,0,w_{j},0,\cdots,0,w_{j-1},0,\cdots,0,w_{1},0,\cdots,0)\in\overline{W_{0}}\subset W^{n},$
for some $0\leq j\leq n$, where $\{x_{1},\cdots,x_{j}\}$ is a linearly independent set. By the Witt extension theorem, we can choose more restrictive representatives of the $(\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F))$-orbits of $\overline{W_{0}}$ as
$(0,\cdots,0,f_{j}^{*},0,\cdots,0,f_{j-1}^{*},0,\cdots,0,f_{1}^{*},0,\cdots,0)\in\overline{W_{0}}\subset W^{n}.$
(Here, $10\leq j\leq n-$ 1 and we set $f_{0}^{*}=0$.) Therefore, there are finite $(\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F))$-orbits in $\overline{W_{0}}$ of the forms
$[(0,\cdots,0,f_{j}^{*},0,\cdots,0,f_{j-1}^{*},0,\cdots,0,f_{1}^{*},0,\cdots,0)]\subset W^{n}$
and index them by $\{\overline{W_{0}}(i)\}_{1\leq i\leq N}$ so that $\dim(\overline{W_{0}}(i))\leq\dim(\overline{W_{0}}(j))$ for $i\leq j$.
Note that for each $j\geq 1$, $\overline{W_{0}}(j)$ is a closed subset of $\bigcup_{i\geq j}\overline{W_{0}}(i)$ and therefore, we have the exact sequence
[…]
We claim that the Schwartz space on each orbits $\overline{W_{0}}$ whose representative is of the form
$(0,\cdots,0,f_{k}^{*},0,\cdots,0,f_{k-1}^{*},0,\cdots,0,f_{1}^{*},0,\cdots,0),\quad\text{for $k<n$}$
is zero.
Let $\overline{W_{0}}(j)$ be an orbit in $\overline{W_{0}}$ whose representative is $\bar{\mathbf{w}}=(0,\cdots,0,f_{k}^{*},0,\cdots,0,f_{k-1}^{*},0,\cdots,0,f_{1}^{*},0,\cdots,0)$ for some $k<n$. Suppose that $\mathcal{S}(\overline{W_{0}}(j))$ is non-zero and put $R_{\bar{\mathbf{w}}}$ the stabilizer of $\bar{\mathbf{w}}$ in $\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$. Consider a map
$\Phi_{\bar{\textbf{w}}}:\mathcal{S}(\overline{W_{0}}(j))\to\operatorname{ind}_{R_{\bar{\mathbf{v}}}}^{\mathrm{Z}_{n}\times\widetilde{\mathrm{J}_{n}}}\mathbb{I},\ \varphi\mapsto\Phi_{\bar{\textbf{w}}}(\varphi),$
where $\Phi_{\bar{\textbf{w}}}$ is defined by
$\Phi_{\bar{\textbf{w}}}(\varphi)(g,h)\coloneqq(\omega_{\psi,V,W}(g,h)\varphi)(\bar{\textbf{w}}).$
It can be readily verified that $\Phi_{\bar{\textbf{w}}}$ is a $(\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F))$-isomorphism. Consequently, we have $\mathcal{S}(\overline{W_{0}}(j))\simeq\text{ind}_{R_{\bar{\mathbf{w}}}}^{\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)}\mathbb{I}$ as a $(\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F))$-module. Since $k<n$, there is a simple root subgroup $J$ of $\mathrm{Z}_{n}(F)$ such that $J\times 1$ is a subgroup of $R_{\bar{\mathbf{w}}}$. However, $\mu$ is non-trivial on $J$ and it leads to a contradiction.
Therefore, by applying the exact sequence (4.7) repeatedly, we have
$\mathcal{S}(W^{n}\oplus X_{n})\simeq\mathcal{S}(\overline{W_{0}})\otimes\mathcal{S}(X_{n})\simeq\mathcal{S}(\overline{W_{0}}(N))\otimes\mathcal{S}(X_{n}),$
where $\overline{W_{0}}(N)$ is a $(\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F))$-orbit in $\overline{W_{0}}$ whose representative is $\bar{\mathbf{w}}^{\prime}=(f_{n}^{*},\cdots,f_{1}^{*})$.
By (4.3), $J_{\mathrm{U},\mu^{-1}}\big{(}\mathcal{S}(\overline{W_{0}}(N))\otimes\mathcal{S}(X_{n}\backslash\{f_{n}\})\big{)}=0$. Therefore, using a similar argument as before, we have
$J_{\mathrm{U},\mu^{-1}}\big{(}\mathcal{S}(W^{n}\oplus X_{n})\big{)}=J_{\mathrm{U},\mu^{-1}}\big{(}\mathcal{S}(\overline{W_{0}}(N))\otimes\mathcal{S}(\{f_{n}\})\big{)}=\mathcal{S}(\overline{W_{0}}(N))\otimes\mathcal{S}(\{f_{n}\}).$
Let $R$ be the stabilizer of $\bar{\mathbf{w}}^{\prime}$ in $\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)$. Then,
$R=\Biggr{\{}\Biggl{(}z,\begin{pmatrix}z&x\\ &z^{*}\end{pmatrix}\Biggr{)}\in\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)\Biggl{\}},$
where we described the elements of $\widetilde{\mathrm{J}_{n}}$ using the basis $\{f_{1},\cdots,f_{n},f_{n}^{*},\cdots,f_{1}^{*}\}$.
[…]
$\mathcal{S}(\overline{W_{0}}(N))\otimes\mathcal{S}(\{f_{n}\})\simeq\operatorname{ind}_{R}^{\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)}(\mathbb{I}\otimes\psi_{n}),$
where $\psi_{n}:R\to\mathbb{C}^{\times}$ is defined by
$\psi_{n}\Biggl{(}\Biggl{(}z,\begin{pmatrix}z&x\\ &z^{*}\end{pmatrix}\Biggr{)}\Biggr{)}=\psi(\frac{x_{n,1}}{2}).$
Therefore, for $\phi\in\operatorname{ind}_{R}^{\mathrm{Z}_{n}(F)\times\widetilde{\mathrm{J}_{n}}(F)}(\mathbb{I}\otimes\psi_{n})$ and $u^{\prime}=\begin{pmatrix}z&x\\ &z^{*}\end{pmatrix}\in\mathrm{U}^{\prime}(F)$ we have
$\phi\big{(}(1,u^{\prime}\cdot h)\big{)}=\psi(z_{1,2}+\cdots+z_{n-1,n}+\frac{x_{n,1}}{2})\phi((1,h))=\mu_{1}^{\prime}(u^{\prime})\phi\big{(}(1,h)\big{)}\quad\text{for all }h\in\widetilde{\mathrm{J}_{n}}(F),$
where $1$ is the identity element in $\mathrm{Z}_{n}(F)$. This proves our claim.
5. The proof
In this section, we will prove Theorem 2.5. To do so, we will make use of the Local Converse Theorem (LCT) for $\mathrm{H}_{n}$. Before proceeding, we will recall a result by Jo that improves upon the earlier work of Jiang and Sourdry ([JS03]).
Theorem 5.1 ([Jo, Theorem 4.15]).
Let $\pi$ and $\pi^{\prime}$ be irreducible admissible generic representations of $\mathrm{H}_{n}(F)$ such that
$\gamma(s,\pi\times\sigma,\psi)=\gamma(s,\pi^{\prime}\times\sigma,\psi)$
holds for any irreducible generic supercuspidal representation $\sigma$ of $\mathrm{GL}_{i}(F)$ with $1\leq i\leq n$. Then
$\pi\simeq\pi^{\prime}.$
Remark 5.2.
In [Jo, Theorem 4.15], an assumption is made regarding the equality of central characters between $\pi$ and $\pi^{\prime}$. Nonetheless, considering that the center of $\mathrm{H}_{n}(F)$ is the trivial group, the central characters of $\pi$ and $\pi^{\prime}$ are inherently trivial. Therefore, we can safely omit this assumption.
The proof of Theorem 5.3 consists of two steps, with the first step focusing on the tempered cases.
Theorem 5.3.
Let $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ be irreducible admissible $\mu_{1}^{\prime}$-generic tempered representations of $\widetilde{\mathrm{J}_{n}}(F)$ such that for any irreducible supercuspidal representation $\sigma$ of $\mathrm{GL}_{i}(F)$ with $1\leq i\leq n$, the equality
$\gamma(s,\widetilde{\pi}\times\sigma,\psi)=\gamma(s,\widetilde{\pi}^{\prime}\times\sigma,\psi)$
holds. Then we have
$\widetilde{\pi}\simeq\widetilde{\pi}^{\prime}.$
Proof.
We begin by considering the bijection $\Theta_{\psi}$ and applying it to $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$. Since both $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ are $\mu_{1}^{\prime}$-generic and tempered, according to Theorem 3.3, the representations $\Theta_{\psi}(\widetilde{\pi})$ and $\Theta_{\psi}(\widetilde{\pi}^{\prime})$ are irreducible generic tempered representations of $\mathrm{H}_{n}^{+}(F)$. Furthermore, by Theorem 3.7, we have
$\gamma(s,\Theta_{\psi}(\widetilde{\pi})\times\sigma,\psi)=\gamma(s,\widetilde{\pi}\times\sigma,\psi)=\gamma(s,\widetilde{\pi}^{\prime}\times\sigma,\psi)=\gamma(s,\Theta_{\psi}(\widetilde{\pi}^{\prime})\times\sigma,\psi).$
for all irreducible supercuspidal representations $\sigma$ of $\mathrm{GL}_{i}(F)$.
Applying Theorem 5.1, we conclude that $\Theta_{\psi}(\widetilde{\pi})\simeq\Theta_{\psi}(\widetilde{\pi}^{\prime})$. Since $\Theta_{\psi}$ is a bijective correspondence, this implies that $\widetilde{\pi}\simeq\widetilde{\pi}^{\prime}$. ∎
Consider a unitary induced representation
$\operatorname{Ind}_{\widetilde{Q}}^{\widetilde{\mathrm{Sp}}(W)}((\tau_{1}|\det|_{F}^{s_{1}})_{\psi}\otimes\cdots\otimes(\tau_{r}|\det|_{F}^{s_{r}})_{\psi}\otimes\widetilde{\pi}_{0}),$
where the following conditions hold:
  • $W=W_{2n}$ and $W_{0}=W_{2n_{0}}$ are symplectic spaces of dimension $2n$ and $2n_{0}$, respectively.
  • $\widetilde{Q}$ is a parabolic subgroup of $\widetilde{\mathrm{Sp}}(W)$ with the Levi subgroup isomorphic to $\widetilde{\mathrm{GL}}_{n_{1}}\times{{\pm 1}}\cdots\times_{{\pm 1}}\widetilde{\mathrm{GL}}_{n_{r}}\times_{\pm 1}\widetilde{\mathrm{Sp}}(W_{0})$.
  • $\tau_{i}$ is an irreducible unitary supercuspidal representation of $\mathrm{GL}_{n_{i}}(F)$.
  • $s_{i}$ is a real number such that $s_{1}\geq\dots\geq s_{r}>0$.
  • $\widetilde{\pi}_{0}$ is an irreducible $\mu_{1}^{\prime}$-generic tempered representation of $\widetilde{\mathrm{Sp}}(W_{0})(F)$.
An irreducible $\mu_{1}^{\prime}$-generic representation of $\widetilde{\mathrm{J}_{n}}(F)$ is a subquotient of such an induced representation (cf. [Jo, Lemma 4.11]). In this case, we say that it has support $(\tau_{1},\cdots,\tau_{r};\widetilde{\pi_{0}})$ and exponents $(s_{1},\cdots,s_{r})$. It should be noted that this notion of support is different from the usual supercuspidal support because $\widetilde{\pi}_{0}$ is not supercuspidal but tempered.
The following lemma is an extended version of [JS03, Proposition 3.2] and [Jo, Lemma 4.11] from supercuspidal to tempered representations.
Lemma 5.4.
For an irreducible generic tempered representation $\tau$ of $\mathrm{H}_{n}(F)$ and unitary supercuspidal representation $\sigma$ of $\mathrm{GL}_{k}(F)$, the followings hold.
  1. $(\mathrm{i})$
    $\gamma(s,\tau\times\sigma,\psi)$ has no zero for $\operatorname{Re}(s)>0$.
  2. $(\mathrm{ii})$
    If $\gamma(s,\tau\times\sigma,\psi)$ has a real pole at $s=s_{0}$, then the pole must be a simple pole at $s_{0}=1$ and $\sigma\simeq\sigma^{\vee}$.
Proof.
From the definition of local $L$-function (2.1), the zero of $\gamma(s,\tau\times\sigma,\psi)$ comes from the poles of $L(s,\tau\times\sigma)$. However, Proposition 2.1 (vii) implies that $L(s,\tau\times\sigma)$ has no pole for $\operatorname{Re}(s)>0$. Therefore $(\mathrm{i})$ of the theorem follows.
In the proof of [Jo, Proposition 4.12 $(\mathrm{iii})$], the requirement that $\pi$ is supercuspidal was imposed in order to apply the ‘Casselman–Shahidi’ Lemma ([Jo, Lemma 4.10]). However, it has been established in [Kim, Proposition 12.3] that the ‘Casselman–Shahidi Lemma’ also holds for tempered representations. Therefore, except for this specific assumption, all other arguments presented in [Jo, Proposition 4.12 $(\mathrm{iii})$] can be equally applied to tempered representations. Hence, property $(\mathrm{ii})$ is proven. ∎
Proposition 5.5.
Let $\sigma$ be an irreducible unitary supercuspidal representation of ${\mathrm{GL}}_{k}(F)$ and $\tau_{i}$, $s_{i}$ and $\widetilde{\pi_{0}}$ are as in the above. Then we have the followings;
  1. $(\mathrm{i})$
    If the product $\prod_{i=1}^{t}\gamma(s+z_{i},\tau_{i}\times\sigma,\psi)$ has a real pole ((((respectively, a real zero)))) at $s=s_{0}$, then $\sigma\cong\tau_{i}^{\vee}$ and $s_{0}=1-z_{i}$ ((((respectively, $s_{0}=-z_{i})$ for some $1\leq i\leq t$.
  2. $(\mathrm{ii})$
    If the product $\prod_{i=1}^{t}\gamma(s-z_{i},\tau_{i}^{\vee}\times\sigma,\psi)$ has a real pole ((((respectively, a real zero)))) at $s=s_{0}$, then $\sigma\cong\tau_{i}$ and $s_{0}=1+z_{i}$ ((((respectively, $s_{0}=z_{i})$ for some $1\leq i\leq t$.
  3. $(\mathrm{iii})$
    The factor $\gamma(s,\widetilde{\pi}_{0}\times\sigma,\psi)$ has no zero for ${\operatorname{Re}}(s)>0$. If $\gamma(s,\widetilde{\pi_{0}}\times\sigma,\psi)$ has a real pole at $s=s_{0}$, then the pole must be a simple pole at $s_{0}=1$ and $\sigma\cong\sigma^{\vee}$.
Proof.
$(\mathrm{i})$ and $(\mathrm{ii})$ follows from the discussion of [JS03, Section 3.2]. For $(\mathrm{iii})$, consider $\Theta_{\psi}(\widetilde{\pi_{0}})$. Then by Theorem 3.1 and Theorem 3.7, $\Theta_{\psi}(\widetilde{\pi_{0}})$ is tempered and
$\gamma(s,\widetilde{\pi_{0}}\times\sigma,\psi)=\gamma(s,\Theta_{\psi}(\widetilde{\pi_{0}})\times\sigma,\psi).$
Then by Lemma 5.4, $(\mathrm{iii})$ readily follows. ∎
The proposition stated below is a consequence of Proposition 5.5. Although the proof is similar to that of [JS03, Theorem 5.1], due to its importance, we provide the proof.
Proposition 5.6.
Let $\widetilde{\pi}$ (resp. $\widetilde{\pi}^{\prime}$) be an irreducible $\mu_{1}^{\prime}$-generic representations of $\widetilde{\mathrm{J}_{n}}(F)$, which has support $(\tau_{1},\cdots,\tau_{r};\widetilde{\pi_{0}})$ (resp. $(\tau_{1}^{\prime},\cdots,\tau_{r^{\prime}};\widetilde{\pi_{0}}^{\prime})$) and exponents $(s_{1},\cdots,s_{r})$ (resp. $(s_{1}^{\prime},\cdots,s_{r^{\prime}}^{\prime})$). Suppose that
(5.1) $\left[\prod_{i=1}^{r}\gamma(s+s_{i},\tau_{i}\times\sigma,\psi)\gamma(s-s_{i},\tau_{i}^{\vee}\times\sigma,\psi)\right]\gamma(s,\widetilde{\pi_{0}}\times\sigma,\psi)\\ =\left[\prod_{i=1}^{r^{\prime}}\gamma(s+s_{i}^{\prime},\tau_{i}^{\prime}\times\sigma,\psi)\gamma(s-s_{i}^{\prime},(\tau_{i}^{\prime})^{\vee}\times\sigma,\psi)\right]\gamma(s,\widetilde{\pi_{0}}^{\prime}\times\sigma,\psi)$
for all irreducible unitary supercuspidal representations $\sigma$ of ${\mathrm{GL}}_{t}(F)$ with $1\leq t\leq n$. Then $r=r^{\prime}$ and there exists a permutation $\rho$ of $\{1,2,\dotsm,r\}$ such that
  1. $(\mathrm{i})$
    $s_{i}=s^{\prime}_{\rho(i)}$ and $\tau_{i}\cong\tau_{\rho(i)}^{\prime}$ for all $i=1,2,\dotsm,r$;
  2. $(\mathrm{ii})$
    $\gamma(s,\widetilde{\pi_{0}}\times\sigma,\psi)=\gamma(s,\widetilde{\pi_{0}}^{\prime}\times\sigma,\psi)$ for all irreducible unitary supercuspidal representations $\sigma$ of ${\mathrm{GL}}_{t}(F)$ with $1\leq t\leq n$.
Proof.
Since the equality (5.1) holds for any irreducible supercuspidal representation $\rho$ of $\mathrm{GL}_{t}(F)$, put $\rho=\rho_{1}$. Then, we have
(5.2) $\left[\prod_{i=1}^{r}\gamma(s+s_{i},\rho_{i}\times\rho_{1},\psi)\cdot\gamma(s-s_{i},\rho_{i}^{\vee}\times\rho_{1},\psi)\right]\cdot\gamma(s,\widetilde{\pi_{01}}\times\rho_{1},\psi)\\ =\left[\prod_{i=1}^{r^{\prime}}\gamma(s+s_{i}^{\prime},\rho_{i}^{\prime}\times\rho_{1},\psi)\cdot\gamma(s-s_{i}^{\prime},(\rho_{i}^{\prime})^{\vee}\times\rho_{1},\psi)\right]\cdot\gamma(s,\widetilde{\pi_{02}}\times\rho_{1},\psi).$
By Proposition 5.4, the left hand side of (5.2) has a pole at $s=1+s_{1}$. The poles of R.H.S (i.e. right hand side) of (5.2) arise from those of $\prod_{1\leq i\leq r^{\prime}}\gamma(s+s_{i}^{\prime},\rho_{i}^{\prime}\times\rho_{1},\psi)$ or $\prod_{1\leq i\leq r^{\prime}}\gamma(s-s_{i}^{\prime},(\rho_{i}^{\prime})^{\vee}\times\rho_{1},\psi)$ or $\gamma(s,\widetilde{\pi_{02}}\times\rho_{1},\psi)$.
If the pole $s=1+s_{1}$ of R.H.S arises from that of $\prod_{1\leq i\leq r^{\prime}}\gamma(s-s_{i}^{\prime},(\rho_{i}^{\prime})^{\vee}\times\rho_{1},\psi)$, then by Proposition 5.4 (ii), $1+s_{1}=1+s_{i}^{\prime}$ and $\rho_{1}=\rho_{i}^{\prime}$ for some $1\leq i\leq r^{\prime}$. Therefore, we can cancel out $\gamma(s+s_{1},\rho_{1}\times\rho,\psi)\cdot\gamma(s-s_{1},\rho_{1}^{\vee}\times\rho,\psi)$ and $\gamma(s+s_{i}^{\prime},\rho_{i}^{\prime}\times\rho,\psi)\cdot\gamma(s-s_{i}^{\prime},(\rho_{i}^{\prime})^{\vee}\times\rho,\psi)$ on both sides of the equality (LABEL:GD). We can perform a similar task for $s=1+s_{2}$. By repeating this process, there is some $1\leq t\leq r$ and a permutation $\mathfrak{p}$ of $\{1,2,\cdots,r^{\prime}\}$ such that $s=1+s_{t}$ is not a pole of $\prod_{i=1}^{r^{\prime}-(t-1)}\gamma(s-s_{\mathfrak{p}(i)}^{\prime},(\rho_{\mathfrak{p}(i)}^{\prime})^{\vee}\times\rho_{t},\psi)$ and the following
(5.3) $\left[\prod_{i=t}^{r}\gamma(s+s_{i},\rho_{i}\times\rho,\psi)\cdot\gamma(s-s_{i},\rho_{i}^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{01}}\times\rho,\psi)\\ =\left[\prod_{i=1}^{r^{\prime}-(t-1)}\gamma(s+s_{\mathfrak{p}(i)}^{\prime},\rho_{\mathfrak{p}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s-s_{\mathfrak{p}(i)}^{\prime},(\rho_{\mathfrak{p}(i)}^{\prime})^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{02}}\times\rho,\psi)$
holds. Note that $s=1+s_{t}$ is a pole of either $\prod_{i=1}^{r^{\prime}-(t-1)}\gamma(s+s_{\mathfrak{p}(i)}^{\prime},\rho_{\mathfrak{p}(i)}^{\prime}\times\rho_{t},\psi)$ or $\gamma(s,\widetilde{\pi_{02}}\times\rho_{t},\psi)$.
If $s=1+s_{t}$ is a pole of $\gamma(s,\widetilde{\pi_{02}}\times\rho_{t},\psi)$, then by Proposition 5.4 (iv), $s_{t}=0$ and $\rho_{t}=\rho_{t}^{\vee}$. Then $\gamma(s,\rho_{t}\times\rho_{t},\psi)^{2}$ has a double pole at $s=1$, while $\gamma(s,\widetilde{\pi_{02}}\times\rho_{t},\psi)$ has a simple pole at $s=1$. Therefore, from (5.3) with $\rho=\rho_{t}$, we see that there is some $1\leq i\leq r^{\prime}-(t-1)$ such that $\gamma(s+s_{\mathfrak{p}(i)}^{\prime},\rho_{\mathfrak{p}(i)}^{\prime}\times\rho_{t},\psi)$ has a pole at $s=1$. It means that $s_{\mathfrak{p}(i)}^{\prime}=0$ and $\rho_{t}=(\rho_{\mathfrak{p}(i)}^{\prime})^{\vee}$. Therefore, if $s=1+s_{t}$ is a pole of $\gamma(s,\widetilde{\pi_{02}}\times\rho_{t},\psi)$, we can cancel out $\gamma(s+s_{t},\rho_{t}\times\rho,\psi)\cdot\gamma(s-s_{t},\rho_{t}^{\vee}\times\rho,\psi)$ and $\gamma(s+s_{\mathfrak{p}(i)}^{\prime},\rho_{\mathfrak{p}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s-s_{\mathfrak{p}(i)}^{\prime},(\rho_{\mathfrak{p}(i)}^{\prime})^{\vee}\times\rho,\psi)$ on both sides of the equality (5.3).
By repeating this process, we arrive at the following equality
(5.4) $\left[\prod_{i=t_{1}}^{r}\gamma(s+s_{i},\rho_{i}\times\rho,\psi)\cdot\gamma(s-s_{i},\rho_{i}^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{01}}\times\rho,\psi)\\ =\left[\prod_{i=1}^{r^{\prime}-(t_{1}-1)}\gamma(s+s_{\mathfrak{p}_{1}(i)}^{\prime},\rho_{\mathfrak{p}_{1}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s-s_{\mathfrak{p}_{1}(i)}^{\prime},(\rho_{\mathfrak{p}_{1}(i)}^{\prime})^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{02}}\times\rho,\psi)$
for some $t\leq t_{1}\leq r$ and a permutation $\mathfrak{p}_{1}$ of $\{1,2,\cdots,r^{\prime}\}$ such that $s=1+s_{t_{1}}$ is not a pole of $\prod_{i=1}^{r^{\prime}-(t_{1}-1)}\gamma(s-s_{\mathfrak{p}_{1}(i)}^{\prime},(\rho_{\mathfrak{p}_{1}(i)}^{\prime})^{\vee}\times\rho_{t_{1}},\psi)$ and $\gamma(s,\widetilde{\pi_{02}}\times\rho_{t_{1}},\psi)$.
[…]
(5.5) $\left[\prod_{i=t_{1}}^{r}\gamma(s,\rho_{i}\times\rho,\psi)\cdot\gamma(s,\rho_{i}^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{01}}\times\rho,\psi)\\ =\left[\prod_{i=1}^{r^{\prime}-(t_{1}-1)}\gamma(s,\rho_{\mathfrak{p}_{1}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s,(\rho_{\mathfrak{p}_{1}(i)}^{\prime})^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{02}}\times\rho,\psi).$
Since $\rho_{t_{1}}=(\rho_{\mathfrak{p}_{1}(i_{1})}^{\prime})^{\vee}$, we can cancel out $\gamma(s,\rho_{t_{1}}\times\rho,\psi)\cdot\gamma(s,\rho_{t_{1}}^{\vee}\times\rho,\psi)$ and $\gamma(s,\rho_{\mathfrak{p}_{1}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s,(\rho_{\mathfrak{p}_{1}(i)}^{\prime})^{\vee}\times\rho,\psi)$ on both sides of the equality (5.5). In this way, we can cancel out $\left[\prod_{i=t_{1}}^{r}\gamma(s,\rho_{i}\times\rho,\psi)\cdot\gamma(s,\rho_{i}^{\vee}\times\rho,\psi)\right]$ in $\left[\prod_{i=1}^{r^{\prime}-(t_{1}-1)}\gamma(s,\rho_{\mathfrak{p}_{1}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s,(\rho_{\mathfrak{p}_{1}(i)}^{\prime})^{\vee}\times\rho,\psi)\right]$ and in turn, we have
(5.6) $\gamma(s,\widetilde{\pi_{01}}\times\rho,\psi)=\left[\prod_{i=1}^{r^{\prime}-r}\gamma(s,\rho_{\mathfrak{p}_{2}(i)}^{\prime}\times\rho,\psi)\cdot\gamma(s,(\rho_{\mathfrak{p}_{2}(i)}^{\prime})^{\vee}\times\rho,\psi)\right]\cdot\gamma(s,\widetilde{\pi_{02}}\times\rho,\psi)$
for some permutation $\mathfrak{p}_{2}$ of $\{1,2,\cdots,r^{\prime}\}$. If $r^{\prime}>r$, put $\rho=\rho_{\mathfrak{p}_{2}(1)}^{\prime}$ in (5.6). Then $\gamma(s,(\rho_{\mathfrak{p}_{2}(1)}^{\prime})^{\vee}\times\rho_{\mathfrak{p}_{2}(1)}^{\prime},\psi)$ has a pole at $s=1$ and so $\gamma(s,\widetilde{\pi_{01}}\times\rho_{\mathfrak{p}_{2}(1)}^{\prime},\psi)$ should have a pole at $s=1$, which implies $\rho_{\mathfrak{p}_{2}(1)}^{\prime}=(\rho_{\mathfrak{p}_{2}(1)}^{\prime})^{\vee}$. However, while the R.H.S of (5.6) with $\rho=\rho_{\mathfrak{p}_{2}(1)}^{\prime}$ has at least double pole at $s=1$, $\gamma(s,\widetilde{\pi_{01}}\times\rho_{\mathfrak{p}_{2}(1)}^{\prime},\psi)$ has at most a simple pole at $s=1$. This is a contradiction. Therefore, we deduce that $r^{\prime}=r$ and it completes the proof. ∎
We are now prepared to prove Theorem 2.5.
Proof.
Assume that $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ have supports as stated in Proposition 5.6. By Proposition 2.1 (i), Theorem 5.3, and Proposition 5.6, we conclude that $\widetilde{\pi}_{0}\simeq\widetilde{\pi}_{0}^{\prime}$, which implies that $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ have the same supports. Since an induced representation has a unique generic constituent, we have $\widetilde{\pi}\simeq\widetilde{\pi}^{\prime}$. ∎
Using similar arguments employed thus far, we can establish Theorem 1.2 as follows.
Proof.
Assume that $\widetilde{\pi}$ has support $(\tau_{1},\cdots,\tau_{r};\widetilde{\pi}_{0})$ and exponents $(s_{1},\cdots,s_{r})$, and $\widetilde{\pi}^{\prime}$ has support $(\tau_{1}^{\prime},\cdots,\tau_{r^{\prime}}^{\prime};\widetilde{\pi}_{0}^{\prime})$ and exponents $(s_{1}^{\prime},\cdots,s_{r^{\prime}}^{\prime})$. Consider the irreducible generic representation $\pi$ of $\mathrm{H}_{n}^{+}(F)$, which has support $(\tau_{1},\cdots,\tau_{r};\Theta_{\psi}(\widetilde{\pi}_{0}))$ and exponents $(s_{1},\cdots,s_{r})$, and $\pi^{\prime}$ of $\mathrm{H}_{n}^{+}(F)$, which has support $(\tau_{1}^{\prime},\cdots,\tau_{r^{\prime}}^{\prime};\Theta_{\psi}(\widetilde{\pi}_{0}^{\prime}))$ and exponents $(s_{1}^{\prime},\cdots,s_{r^{\prime}}^{\prime})$.
By [CKPSS04, Theorem 4.1], [CP, Theorem], and [L15, Theorem 1.4], there exists an integer $l_{0}=l(\pi,\pi^{\prime})>0$ such that for any quasi-character $\chi$ of $F^{\times}$ with $\text{cond}(\chi)>l_{0}$,
$\gamma(s,\pi\times\chi,\psi)=\gamma(s,\pi^{\prime}\times\chi,\psi).$
Furthermore, by Theorem 3.7, we know that
$\gamma(s,\widetilde{\pi_{0}}\times\chi,\psi)=\gamma(s,\Theta_{\psi}(\widetilde{\pi_{0}})\times\chi,\psi),\quad\gamma(s,\widetilde{\pi_{0}}^{\prime}\times\chi,\psi)=\gamma(s,\Theta_{\psi}(\widetilde{\pi_{0}}^{\prime})\times\chi,\psi).$
Therefore,
$\gamma(s,\widetilde{\pi}\times\chi,\psi)=\gamma(s,\pi\times\chi,\psi),\quad\gamma(s,\widetilde{\pi}^{\prime}\times\chi,\psi)=\gamma(s,\pi^{\prime}\times\chi,\psi).$
By taking $l=l_{0}$, this completes the proof. ∎
Remark 5.7.
Indeed, the proof of Theorem 1.2 does not essentially require tempered supports (i.e., $\widetilde{\pi}_{0}$ and $\widetilde{\pi}_{0}^{\prime}$ being tempered), but supercuspidal support is sufficient.
6. An application
We present a direct application of our main theorem, which is the rigidity theorem for metaplectic groups. The discussion in this section is similar to [JS03, Theorem 5.3].
In this section, let $K$ be a number field and $\mathbb{A}$ its adéle ring. Let $\mathfrak{U}^{\prime}$ be a generic character of $\mathrm{U}^{\prime}(K)\backslash\mathrm{U}^{\prime}(\mathbb{A})$. Let $\widetilde{\pi}=\otimes_{v}\widetilde{\pi}_{v}$ and $\Pi=\otimes_{v}\Pi_{v}$ be irreducible automorphic representations of $\widetilde{\mathrm{Sp}}_{2n}(\mathbb{A})$ and $\mathrm{GL}_{2n}(\mathbb{A})$, respectively. Then we say that $\Pi$ is a weak functorial lift of $\widetilde{\pi}$ if $\Pi_{v}$ is the local Langlands functorial lift of $\widetilde{\pi_{v}}$ for all archimedean places and for almost all places $v$ of $F$, where $\widetilde{\pi_{v}}$ and $\Pi_{v}$ are unramified.
Lemma 6.1.
Let $\widetilde{\pi}=\otimes_{v}\widetilde{\pi_{v}}$ be an irreducible $\mathfrak{U}^{\prime}$-generic cuspidal automorphic representation of $\widetilde{\mathrm{Sp}}_{2n}(\mathbb{A})$. Let $\Pi=\otimes_{v}\Pi_{v}$ be any weak functorial lift of $\widetilde{\pi}$. Fix any finite place $v$ of $K$. Then for every supercuspidal generic representation $\sigma_{v}$ of $\mathrm{GL}_{t}(F_{v})$ with $1\leq t\leq 2n$, we have
$\gamma(s,\pi_{v}\times\sigma_{v},\psi_{v})=\gamma(s,\Pi_{v}\times\sigma_{v},\psi_{v}).$
Proof.
The proof follows a similar approach to [CKPSS, Corollary 4, pp.26]. The main ingredient in the proof of [CKPSS, Corollary 4] is the stability of local $\gamma$-factors for $\mathrm{SO}_{2n+1}$. However, Theorem 1.2 plays a similar role for metaplectic groups, and the remaining arguments work equally well. ∎
Now we are ready to prove Theorem 1.3.
Proof.
Let $\Pi$ and $\Pi^{\prime}$ be a weak functorial lifts of $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ to $\mathrm{GL}_{2n}(\mathbb{A})$. Then by [GRS, Theorem 11.2], $\Pi$ and $\Pi^{\prime}$ are irreducible generic self-dual unitary automorphic representations of $\mathrm{GL}_{2n}(\mathbb{A})$ and by the assumption, $\Pi_{v}\simeq\Pi_{v}^{\prime}$ for almost all places $v$. Then by the strong multiplicity one theorem for $\mathrm{GL}_{2n}(\mathbb{A})$, $\Pi\simeq\Pi^{\prime}$ and thus $\Pi_{v}\simeq\Pi_{v}^{\prime}$ for all places $v$. Then by the property of the local functorial lifts, $\widetilde{\pi}_{v}\simeq\widetilde{\pi}^{\prime}_{v}$ for all archimedean places. Furthermore, by Lemma 6.1, for arbitrary finite place $v$, we have
$\gamma(s,\widetilde{\pi}_{v}\times\sigma,\psi)=\gamma(s,\Pi_{v}\times\sigma,\psi)=\gamma(s,\Pi_{v}^{\prime}\times\sigma,\psi)=\gamma(s,\widetilde{\pi}^{\prime}_{v}\times\sigma,\psi)$
for any irreducible generic supercuspidal representation $\sigma$ of $\mathrm{GL}_{t}(F_{v})$ with $1\leq t\leq n$. Then by Theorem 1.1, we have $\widetilde{\pi}_{v}\simeq\widetilde{\pi}^{\prime}_{v}$. This completes the proof of the theorem. ∎
Acknowledgements
The author would like to express gratitude to the anonymous referee for their helpful comments, which have improved the exposition of this paper. Special thanks to Hiraku Atobe and Eyal Kaplan for their valuable comments. The author also extends heartfelt thanks to Dongho Byeon, Youngju Choie, Youn-Seo Choi, Wee Teck Gan, Haseo Ki, and Sugwoo Shin for their unwavering support and encouragement over the years.
Finally, the author expresses gratitude to Yonsei University and Catholic Kwandong University for providing an excellent research environment and for their generous support. This work is supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) under grant number 2020R1F1A1A01048645.